### v1

# TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching

Abstract
To guide a learner to master the action skills, it is crucial for a coach to 1) reason through the learner’s action execution and technical keypoints, and 2) provide detailed, understandable feedback on what is done well and what can be improved. However, existing score-based action assessment methods are still far from this practical scenario. To bridge this gap, we investigate a new task termed Descriptive Action Coaching (DAC) which requires a model to provide detailed commentary on what is done well and what can be improved beyond a quality score from an action execution. To this end, we construct a new dataset named EE4D-DAC. With an LLM-based annotation pipeline, our dataset goes beyond the existing action assessment datasets by providing the hierarchical coaching commentary at both keypoint and instance levels. Furthermore, we propose TechCoach, a new framework that explicitly incorporates keypoint-level reasoning into the DAC process. The central to our method lies in the Context-aware Keypoint Reasoner, which enables TechCoach to learn keypoint-related quality representations by querying visual context under the supervision of keypoint-level coaching commentary. Prompted by the visual context and the keypoint-related quality representations, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. Combining all of these, we build a new benchmark for DAC and evaluate the effectiveness of our method through extensive experiments. Data and code will be publicly available.
footnotetext: *: Corresponding author. $\dagger$: Equal contributions.
1 Introduction
Figure 1: An illustration of real-world action coaching. Upper: Given an action demonstration, a human coach will provide coaching comments on what is done well and what can be improved. Lower: To provide precise feedback, the coach incorporates Keypoint-level Reasoning into the action coaching process, i.e., keep the technical keypoints (e.g., The player should extend the arms fully as they bring the ball under the basket) in mind and determine the keypoint-level action quality (e.g., The athlete performs beautifully with the right arm extension) of the action. Such a reasoning process motivates us to propose our TechCoach to address Descriptive Action Coaching.
Understanding how well an action is performed, also known as Action Quality Assessment (AQA), has recently attached growing attention due to its potential applications [38, 45, 29, 8]. One of the most promising potentials of AQA is to build an AI action coach, guiding an action learner toward gradually mastering the skill. While impressive progress has been achieved, current AQA methods are still far from real coaching and action-guiding scenarios, especially in their functionality and reasoning process:
- Functionality. In practical action-guiding scenarios, a coach must provide detailed and understandable feedback on what is done well and what can be improved so that the learner can be fully aware of the execution details and master the skills (see the upper part of Fig. 1). However, existing works pose AQA as a score-regression [28, 35, 50] or pairwise-ranking [9, 55, 12] problem, limiting the applicability of the models to more practical scenarios.
- Reasoning process. To provide precise feedback, a coach will keep the technical keypoints (e.g., The player should extend the arms fully in Reverse Layup) in mind and determine the keypoint-level action quality from action execution (see the lower part of Fig. 1). However, few existing works have incorporated such a keypoint-level reasoning process into AQA. While recent works [39, 15] propose to address this issue by recognizing the existence of keypoint-level mistakes, such a paradigm still limits deeper exploration of the connections between the action execution and the keypoint-level action quality (i.e., detailed strength/weakness on a technical keypoint).
To bridge the gap between the field of AQA and real-world action coaching scenarios, we investigate a new task termed Descriptive Action Coaching (DAC), which requires the model to provide detailed commentary on what is done well and what can be improved, beyond merely assigning a quality score to an action execution. Such a task is challenging because to provide such a detailed commentary, the model not only needs to understand fine-grained action execution (e.g., the player looks down when performing Reverse Layup), but also build explicit connections between the execution with the action quality (e.g., it is a weakness to look down at the ball as it will miss the target).
To facilitate the studies on DAC, we construct EE4D-DAC, a new dataset that not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint and instance levels. Specifically, we first source the action videos from the recently proposed EgoExo4D [10] dataset. Subsequently, an LLM-driven automatic annotation pipeline is proposed to progressively obtain the general technical keypoints and hierarchical coaching commentary.
The construction of the EE4D-DAC dataset allows us to build Technical Keypoint-aware Descriptive Action Coach (TechCoach), a new DAC framework which explicitly incorporates keypoint-level reasoning into the action coaching process. The key idea of TechCoach is to imitate the real-world coaching scenario, where the coach first reasons the action quality on each technical keypoint, and then provides feedback by incorporating the keypoint-level reasoning results and visual context.
To this end, TechCoach includes a Context-aware Keypoint Reasoner, which queries the visual context with general technical keypoints to obtain keypoint-related quality representations. To ensure such representations carry the action quality information, a keypoint-level alignment loss is proposed to align the representations with the keypoint-level coaching commentary. After that, a Unified Keypoint-aware Action Assessor (KA2) is utilized to regress the action quality score and generate overall coaching commentary. In this process, a progressive action coaching attention mask is adopted to guide KA2 to integrate information starting with the visual context level, followed by the keypoint level, and ultimately the decision (i.e., score-prediction and commentary-generation) level.
Combining the EE4D-DAC dataset and TechCoach, we can construct a new benchmark to expand the field of AQA. Extensive experiments demonstrate that: 1) Our TechCoach achieves state-of-the-art performance among the compared methods, including task-specific models and general MLLMs; 2) The Context-aware Keypoint Reasoner is a core design and cannot be simply replaced by other solutions (e.g., replacing the Keypoint-level Alignment with a classification-based solution that classify whether the execution shows strength or weakness on each keypoint).
In summary, the contributions of this work can be summarized as: 1) We construct a new dataset termed EE4D-DAC to address Descriptive Action Coaching. To our knowledge, EE4D-DAC is the first dataset that includes hierarchical coaching commentary on both keypoint level and instance level. 2) We develop TechCoach, a new method to address DAC by explicitly collaborating keypoint-level reasoning into the action coaching process. 3) We construct a new benchmark on DAC. Experiments not only evaluate the effectiveness of the proposed approach but also provide a starting point for future work in this field.
Figure 2: An overview of the LLM-driven automatic annotation pipeline for the EE4D-DAC dataset. We first prompt the LLM to generate general technical keypoints for the given action task (Orange $\rightarrow$). Subsequently, we ask the LLM to summarize the keypoint-level and instance-level commentary according to the general technical keypoints and the original expert commentary (Blue $\rightarrow$). Information about various keypoints is highlighted in colors. Zoom in for best view.
2 Related Works
Action Quality Assessment (AQA) [43, 35, 50, 24, 47, 45, 3, 44, 29, 51, 27, 54, 30, 15, 22, 21, 57] aims at evaluating how well an action is performed. Most existing works pose AQA as a score regression task, and approach this task in many ways, e.g., introducing new formulations [43, 35, 50], augmenting human-body representations [24, 47], exploring temporal structures[45, 3, 44], collaborating multi-modal information [29, 51, 42]. Although impressive progress has been achieved, regressing a single score is still far from real-world action coaching scenarios. While [27, 54] propose to generate commentary in addition to the score, the commentary comes from the live comment, focusing on action and score rather than skill guiding and coaching. Recently, considering it is crucial for a coaching model to reason through various technical keypoints, [30, 15, 39] propose to detect mistakes on various body parts or technical keypoints. [22, 21] introduce a scoring rubric to inject judging rules into the score prediction process. However, This branch of methods is still limited by the binary or score-only feedback. Notably, [23] incorporates multiple models together with judging rules to provide formatted feedback for diving judging, but requires complex manual design on a single task. Differently, our work focuses on generating free-from commentary on what is done well and what can be improved, which is more aligned with practical coaching scenarios. To achieve this, we construct a new dataset that features detailed coaching commentary on both keypoint and instance levels. Furthermore, we propose a new general framework that collaborates key-point reasoning into coaching process.
Video Captioning [34, 46, 17, 40, 19, 49] is a long-lasted task that requires a model to generate language caption with video frames as input. Early studies focus on designing stronger architectures [17, 40, 1, 33, 48] or utilizing video-language pretraining [34, 19, 49, 2, 32] to model relationships between video and language. Recently, Multimodal Large Language Models (MLLMs) [16, 41, 14, 20, 13, 7] show strong generalization ability across various video-language tasks such as video content description and video question answering. In this paper, we focus on Descriptive Action Coaching, a rarely explored but challenging scenario that requires video-language models to collaborate fine-grained video content and domain knowledge to provide detailed coaching commentary. We provide a new testbed for today’s video-language models and show that they still struggle to handle such a practical scenario. Besides, we propose a new keypoint-guided coaching model and evaluate its effectiveness through extensive experiments.
3 EE4D-DAC Dataset
To enable studies on Descriptive Action Coaching, we construct a new dataset termed EE4D-DAC. Different from existing action assessment datasets, EE4D-DAC not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint level and instance level. We will introduce the data source in Sec. 3.1, the annotation pipeline in Sec. 3.2, and provide comparisons with existing datasets in Sec. 3.3 respectively.
3.1 Data Source
We source data from the recently proposed EgoExo4D [10] dataset. Except for various ego-exo action videos, EgoExo4D also contains execution rating (from 1 to 10) and time-aligned free-form spoken commentary (e.g. “Yeah, see how she’s having to look down. You can’t see the basket here.") by domain experts. Such characteristics make the EgoExo4D dataset a great starting point for studying action coaching.
In our work, we focus on physical actions and select takes from EgoExo4D including the scenarios of basketball, soccer, and bouldering. After that, we segment each take into single instances with a fixed-size (8-second) window and filter those instances without meaningful commentary. By doing so, we obtain 4843 unique video instances, spanning about 10.8 hours.
Datasets Years Data Unique Unique KeyPoint Level Judgment Instance Level Judgment
Sources Instances Hours Exists Type Num Score Commentary Coach-Score
Datasets w/o Language Annotations
MIT-Skate [31] ECCV14 Olympics 150 7.3h $\times$ - - $\times$ -
AQA-7 [26] WACV19 Olympics 1189 1.62h $\times$ - - $\times$ -
Rhythmic Gymnastics [52] ACMMM20 Olympics 1000 23.4h $\times$ - - $\times$ -
FineDiving [45] CVPR22 Olympics 3000 3.5h $\times$ - - $\times$ -
LOGO [53] CVPR23 Olympics 200 11.4h $\times$ - - $\times$ -
Skate-IRIS [22] ACMIUI23 MIT-Skate [31] 150 7.3h Score 1050 $\times$ -
CPR-Coach [39] CVPR24 Live Saving 1416 7.7h Binary 18.4k $\times$ $\times$ -
Datasets w/ Language Annotations
MTL-AQA [27] CVPR19 Olympics 1412 1.7h $\times$ - - 1.94
MTL-NAE [54] CVPR24 MTL-AQA [27] 1412 1.7h $\times$ - - 2.94
EgoExo-Fitness [15] ECCV24 Daily Fitness 913 4.6h Binary 7.8k 2.37
EgoExo4D (Sports) [10] CVPR24 Daily Sports 1219 16.4h $\times$ - - 3.44
EE4D-DAC(Ours) EgoExo4D [10] 4843 10.8h Commentary 25.1k 4.33
Table 1: Comparison to popular Action Assessment datasets. Our EE4D-DAC dataset is the first dataset that contains detailed coaching commentary on both key-point level and instance level. Coach-Score: A metric that evaluates whether the provided commentary is suited for real-world coaching scenarios (See Sec. 3.3 for details).
3.2 Annotation Pipeline
By going deeper into the expert commentary in EgoExo4D, it can be observed that: 1) The original expert commentary is highly colloquial and noisy, awaiting summarization before being used to build a coaching model. 2) The expert commentary is highly associated with the general technical keypoints, making it possible to mine the relationships between the general technical keypoints and final coaching commentary.
Based on the observations, we propose an LLM-based automatic annotation pipeline to obtain general technical keypoints, and hierarchical coaching commentary on both keypoint level and instance level, which is shown in Fig. 2. We will bring more details as follows.
- Collection of the General Technical Keypoints. In the general coaching-learning scenario, a coach will provide several technical keypoints in advance so that the skill learner can learn to follow. Considering that completing a skilled physical action always requires cooperation across multiple body parts and objects, we define a technical keypoint as a language instruction on one of the seven dimensions, including six body parts (e.g., head & eyes, arms & hands) and the human-objects interaction.
To obtain technical keypoints for each skilled action task, we treat the LLM (GPT-4O) as a general coach. As shown in the lower left part of Fig. 2, by providing the task name (e.g., Basketball Drills - Reverse Layup) and detailed instructions to the LLM, we are able to obtain the general technical keypoints on the previously mentioned seven dimensions (e.g., For the dimension of “head & eyes”, we have the keypoint: “The player should keep the head up and eyes focused on the backboard and rim to accurately judge the distance and angle for the reverse layup.”).
- Collection of Hierarchical Coaching Commentary. Subsequently, we further ask the LLM to mine the relationships between the technical keypoints and origin expert commentary and provide summarized hierarchical (i.e., keypoint level and instance level) coaching commentary about what is done well and what can be improved in the action execution.
As shown in the right part of Fig. 2, in this phase, LLM is asked to finish two tasks: (1) Review the general technical keypoints and the original expert commentary, and then summarize the strength and weakness in the execution corresponding to each keypoint. (2) Review the original expert commentary and the generated strengths and weaknesses, and then provide the overall commentary on the execution.
To sum up, with the annotation pipeline, for each unique instance, we construct rich language annotations including the General Technical Keypoints, Keypoint-level Commentary, and Instance-level Commentary. Moreover, for each instance we also obtain an average rating by averaging all the ratings provided by different experts.
- Discussions. Note that in some cases, an execution could reveal both the strength and weakness aspects on one keypoint, and some may just reveal one of them. Such a characteristic is inherited from EgoExo4D [10].
3.3 Comparison with existing datasets
We compare our dataset with related AQA datasets. As shown in Tab. 1, the detailed keypoint-level commentary allows us to separate our EE4D-DAC dataset from the other datasets [22, 39, 15] with keypoint-level annotations on both type and scale. Furthermore, we conduct an evaluation to show the guiding ability of the instance-level commentary given by different datasets. Specifically, we prompt the ChatGPT to rate the commentary from 0 to 5 as the Coach-Score. As shown in the last column in Tab. 1, commentary MTL-AQA [27] and MTL-NAE [54] is not suit for coaching scenario as the original text mainly focuses on the action and score rather than action guiding and skill improvement.
For more details about the prompts, the pre-defined body parts, the examples and dataset comparison strategies, please refer to the Appendix.
4 TechCoach
4.1 Problem Formulation
Given an action video $v$, the goal of Descriptive Action Coaching is to train a model that takes the video as input and predicts an overall action quality score together with a paragraph of detailed commentary about what is done well and what can be improved.
Figure 3: An overview of the proposed TechCoach. Given an action video, TechCoach first extracts visual context embeddings with a pre-trained visual backbone and a Spatial-Temporal Context Enhancer. After that, a Context-aware Keypoint Reasoner is adopted to learn keypoint-related quality embeddings by querying visual context under the supervision of keypoint-level coaching commentary. Finally, prompted by the visual context and the keypoint-related quality embeddings, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. K.P.: Keypoints. Best viewed in color.
Figure 4: An overview of the Context-aware Keypoint Reasoner. Best viewed in color.
4.2 Overview
An overview of the proposed Technical Keypoint-aware Descriptive Action Coach (TechCoach) is shown in Fig. 3. The key idea of TechCoach is to mimic the reasoning process of human coaches by incorporating keypoint-level reasoning into the action coaching pipeline. Our framework consists of three parts. Firstly, given an input video $v$, a Visual Encoding Module will extract the visual context embedding. After that, a Context-aware Key-point Reasoner is adopted to take the visual context embedding and general technical keypoints as input to perceive the action quality on each keypoints and provide the key-point quality embeddings. Finally, a Unified Keypoint-aware Action Assessor is utilized to collaborate the visual context and keypoint-aware quality embeddings, then predict the final quality score and generate coaching commentary.
Discussions. Note that our model is based on the assumption that the general technical keypoints are available. We want to argue that this assumption is reasonable because: 1) It is aligned with real-world coaching scenarios that a coach is well-informed about the keypoints and able to reason through the keypoints and learner’s executions. 2) Obtaining keypoints is not resource-intensive as we can prompt LLMs (e.g., ChatGPT) to provide the keypoints.
4.3 Visual Encoding
Given a video $v$, we first divide the video into non-overlapping video clips and extract clip-level features with a well-pretrained backbone (e.g., InternVideo2 [41]). After that, inspired by TimeSformer [5], we utilize a Transformer-based Spatial-Temporal Context Enhancer to enhance the spatial-temporal context across the clip-level features. This process produces the visual context embeddings $f_{v}\in R^{T\times H\times W\times D}$ can be obtained, where $T$, $H$, $W$, $D$ indicate the temporal, height, width, and channel size, respectively. The details about the Spatial-Temporal Context Enhancer are presented in Appendix.
4.4 Context-aware Key-point Reasoner
To generate comprehensive coaching commentary for an action execution, an intuitive approach is to directly input the video (and the keypoint) features into the text generator. However, without explicit modeling, such an approach is unable to effectively perceive the relationship between visual context and the keypoints. To address this, TechCoach adopts a Context-aware Key-point Reasoner to learn keypoint-related quality embeddings by querying the video context embeddings under the supervision of the coaching commentary for each technical keypoint. An overview of Context-aware Keypoint Reasoner is shown in Fig. 4.
- Keypoint Queries Construction. Without loss of generality, given the N general technical keypoints corresponding to the input video, we first extract the keypoint features $f_{kp}\in R^{N\times D}$ with a pre-trained text encoder (e.g., DeBERTa [11]) and a linear mapper. After that, we augment $f_{kp}$ by introducing two types of learnable prompts. Specifically, we use a group of keypoint-type prompts $f_{kt}\in R^{N\times D}$ to identify which dimension (e.g., head & eyes, human-object interaction) each keypoint feature belongs to. Besides, we adopt two quality-related prompts $f_{q}\in R^{2\times D}$ to indicate the aspects of strength and weakness. Based on the definitions of the learnable prompts, the augmented keypoint features $f^{*}_{kp}\in R^{N\times 2\times D}$ are computed as:
$f^{*}_{kp}[i,j]=f_{kp}[i]+f_{kt}[i]+f_{q}[j].\vspace{-0.2cm}$ (1)
Such an augmentation allows us to inject the keypoint and quality information into the keypoint queries.
- Keypoint-driven Quality Reasoning. Our goal here is to decouple the visual context features corresponding to the keypoints and quality aspects. To this end, we regard the augmented keypoint features $f^{*}_{kp}$ as the queries and visual context features $f_{v}$ as the keys and values, then utilize layers of multi-head cross attention module [36] to obtain the keypoint-related quality embeddings $f_{kq}\in R^{N\times 2\times D}$. $f_{kq}[i,j]$ indicates the action quality on keypoint $i$ and aspect $j$ ($1$ for the strength and $2$ for the weakness).
To ensure $f_{kq}$ carries the quality information related to the action execution on the technical keypoints, our idea is to align the keypoint-related quality embeddings $f_{kq}$ with the features of the corresponding keypoint-level coaching commentary. To this end, a keypoint-level alignment loss is adopted, which can be written as:
$L_{Align}=\sum^{N}_{i=1}\sum^{2}_{j=1}w_{kc}[i,j]*D(\psi(f_{kq}[i,j]),f_{kc}[i,j]),\vspace{-0.2cm}$ (2)
where $f_{kc}$ indicates the features of keypoint-level commentary extracted by the text encoder; $\psi$ is a linear mapper that projects the $f_{kq}$ back to the dimension of the text features; $w_{kc}[i,j]$ is a binary scalar to filter those cases where commentary on some aspects is not provided (see Discussions in Sec. 3.2); $D(\cdot,\cdot)$ indicates a distance measurement between two features, and we use the L2 Distance by default.
4.5 Unified Keypoint-aware Action Assessor
After obtaining the visual context embeddings and the keypoint-related quality embeddings, we employ a Transformer-based Unified Keypoint-aware Action Assessor (KA2) to jointly predict the action quality score and generate the overall coaching commentary.
Specifically, the input sequence to the KA2 module is constructed by concatenating the following components: (1) The generated text tokens $T_{1:i-1}$; (2) a learnable score-prediction token $f_{s}$; (3) the flattened $f_{kc}$ as keypoint tokens; (4) the flattened $f_{v}$ as video tokens. For generating commentary, we follow [17, 54] by employing a Mask-Token Prediction supervision ($L_{MTM}$) during training and using the Next-Token Prediction paradigm during inference. For score regression, the score-prediction token from the output layer is passed to a fully connected regressor, optimized using a Mean Squared Error (MSE) loss ($L_{MSE}$).
Moreover, to guide the KA2 module in progressively integrating information in a sequence from low-level to high-level (i.e., starting with visual context level, then keypoint level, and finally the decision level), a progressive action coaching attention mask is further adopted on the KA2, which is shown in the lower right part of Fig. 3. Specifically, each keypoint token independently attends to itself and the video tokens. Besides, the score-prediction token attends to itself and all the keypoint and video tokens. For commentary generation, KA2 integrates all information by attending to the generated text tokens, the score-prediction token, all the keypoint tokens, and the video tokens.
To sum up, we train the TechCoach with the following overall loss function:
$L=L_{MTM}+\lambda_{1}L_{MSE}+\lambda_{2}L_{Align},\vspace{-0.2cm}$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters.
Methods Years Score Regression
$\rho\uparrow$ RL2$\downarrow$
Models infer with single instance
InternVideo2-MLP - 69.75 4.70
USDL [35] CVPR21 70.81 4.51
SwinBERT [17] CVPR22 71.37 4.59
PGMI [54] CVPR24 70.93 4.38
TechCoach(Ours) 72.46 4.35
Models infer with extra exemplars
CoRe [50] w/ 10 Exem. ICCV21 73.17 4.35
TPT[3] w/ 10 Exem. ECCV22 73.77 4.29
(a) Comparison with existing methods on score regression.
Methods Years Commentary Generation
[…]
General Multimodal Large Language Models
VideoLLaVA-7B [16] ACL24 22.94 1.79 9.80 53.76 16.62 14.82
VideoChat2-7B [14] CVPR24 23.07 3.20 11.94 60.79 17.82 25.99
InternVideo2-Stage3-8B [41] ECCV24 27.52 2.98 12.83 63.80 37.85 28.86
InternVL2-8B [6] 2024 12.51 0.10 16.81 62.30 47.76 34.01
Task-Specific Models
SwinBERT[17] CVPR22 36.11 11.70 16.01 66.03 46.56 34.57
PGMI [54] CVPR24 36.78 14.00 16.24 66.42 49.02 37.04
TechCoach(Ours) - 37.06 14.62 16.39 66.89 50.44 38.15
(b) Comparison with existing methods on coaching commentary generation.
Table 2: Comparison with existing models. Our TechCoach outperforms both the general MLLMs and Multi-Task Learning-based AQA models on coaching commentary generation. Besides, TechCoach shows strong score regression abilities with the highest performance among the methods using single instance for predicting the score. “w/ 10 Exem.”: We gray out models that infer with 10 extra exemplars for fair comparisons.
5 Experiments
5.1 Dataset and Evaluation Metrics
- Dataset. We conduct experiments on EE4D-DAC to provide a new benchmark and evaluate the performance of the proposed TechCoach. Following the official split of EgoExo4D [10], we separate our EE4D-DAC into training and evaluation set (including 3769 and 1074 instances, respectively).
- Evaluation Metrics. For score prediction, we use the Spearman’s rank correlation coefficient ($\rho$) and Relative-L2 Distance (RL2) as the metrics. For natural language generation, we adopt NLP metrics, including BERT [56], BLEU(B)[25], METEOR(M)[4] and CIDEr(C) [37] scores. Moreover, for better comparison with open-source MLLMs, we design two more LLM-based Metrics: (1) a Mention Score(GPT-M) to evaluate whether the generated commentary mentions the same technical details as in the ground truth; (2) a Quality Score(GPT-Q) to evaluate whether the generated commentary shares the same praises or improvement opinions as in the ground truth on those both-mentioned technical details. For more details about the evaluation metrics, please refer to the Appendix.
5.2 Implementation Details
For each video instance, we sample frames and segment them into 32 8-frame clips. Subsequently, we use pretrained InternVideo2 [41] to extract clip features and adopt average pooling to reduce the spatial and temporal resolutions. obtaining video features with a size of $16(T)\times 8(H)\times 8(W)$. We use the pretrained DeBERTa-xlarge [11] as the text encoder. We use both the ego- and exo-centric (best-exo in EgoExo4D) videos as input. We concatenate the multi-view visual features on channel dimension and feed them into a linear projector before going through the Spatial-Temporal Context Enhancer. For more implementation details, please refer to the Appendix.
5.3 Main Results
We compare the proposed methods with two branches of baselines: (1) Task-Specific Models: We select AQA- and VideoCaption-specific baselines including USDL [35], CoRe [50], TPT[3], SwinBERT[17] and PGMI [54]. For fair comparisons, we evaluate all domain-specific models by using the same visual backbone (i.e., InternVideo2) to extract features and the same fusion strategy on multi-view features. (2) General Multi-modal Large Language Models: We also select popular open-source MLLMs as the baselines, including VideoLLaVA [16], VideoChat2 [14], InternVideo2-Stage3 [41], and InternVL2 [6]. For more details about the baseline models, please refer to the Appendix. The comparison results are shown in Tab. 2 and we have the following observations:
Variants Score Commentary
[…]
70.78 4.45 - - - -
- - 36.71 13.70 16.39 66.77
70.23 4.77 36.42 14.26 16.09 65.72
72.46 4.35 37.06 14.62 16.39 66.89
Table 3: Ablation studies on training losses. The results show that keypoint-level alignment is a core design to achieve a better trade-off between score regression and coaching commentary generation.
Variants Score Commentary
[…]
w/o KP-Align 70.23 4.77 36.42 14.26 16.09 65.72
KP-CLS 72.13 4.45 36.78 13.69 16.19 64.69
KP-Align 72.46 4.35 37.06 14.62 16.39 66.89
Table 4: Ablation studies on the alternative solutions of keypoint-level alignment. The results show that the proposed keypoint-level alignment (KP-Align) cannot be simply replaced by the naive classification-based solution (KP-CLS).
1) Score Regression: Compared with Direct Regression-based (i.e., InternVideo2-MLP and USDL) and Multi-Task Learning-based (i.e., SwinBERT, PGMI) AQA methods, TechCoach achieves the best performance on score regression. Note that CoRe and TPT adopt a Contrastive Regression framework that utilizes 10 extra exemplars during inference, which brings stronger score regression ability.
2) Coaching Commentary Generation: Although recently proposed Multimodal Large Language Models (MLLMs) show great visual context understanding and reasoning abilities, most of them [41, 14, 16] still struggle to handle realistic coaching scenarios (under-perform the task-specific models). Notably, the best among the evaluated MLLMs is the InternVL2 [6], which shows comparable performance with the task-specific models on LLM-based metrics. While InternVL2 is able to mention similar technical details as the ground truth (with a 47.76 GPT-M score), it falls short in providing precise guiding opinions (only 34.01 on the GPT-Q score). Among all the compared methods, our TechCoach achieves the best performance among all the metrics on coaching commentary generation.
Variants Score Commentary
[…]
70.83 4.49 36.25 13.34 15.88 66.36
71.96 4.47 35.99 13.24 15.95 66.51
72.00 4.39 36.43 12.57 16.07 66.72
72.46 4.35 37.06 14.62 16.39 66.89
Table 5: Ablation studies on the impact of Keypoint Features Augmentation. K.P.Text: Keypoint text features. K.P.Type: keypoint-type prompts. Quality: quality-related prompts.
Variants Score Commentary
[…]
Ego-Only 69.31 4.75 36.87 13.79 16.16 66.13
Exo-Only 68.55 4.84 37.31 12.71 16.20 66.23
Ego+Exo 72.46 4.35 37.06 14.62 16.39 66.89
Table 6: Comparison with Ego- and Exo-only variants. By collaborating on both Ego and Exo videos, TechCoach achieves stronger performance, showcasing the complementary role across different views for descriptive action coaching.
Figure 5: Qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors.
5.4 Ablation Studies
- Is keypoint-level alignment a necessary? We study this question by gradually ablating the training losses illustrated in Eq. 3 from our full model. The results are shown in Tab. 3. By comparing Row 3 with Row 1-2, it can be clearly observed that the model achieves unstable performance when simply combining the $L_{MSE}$ and $L_{MTM}$. This may caused by the gap between the score regression and commentary generation tasks. However, when further combining the keypoint-level alignment loss (Row 4), we can see stable performance improvement on all the compared metrics. Hence, we can conclude that keypoint-level alignment is necessary for the trade-off of score regression and coaching commentary generation.
- Does keypoint-level alignment equal to binary classification? The main target of the keypoint-level alignment (KP-Align) is to ensure the Keypoint-related Quality Embeddings (see Fig. 3 and Sec. 4.4 for more details) carry the quality-related information. To achieve this, a simple solution is to perform binary classifications over the Keypoint-related Quality Embeddings (i.e., classify whether the execution shows strengths or weaknesses on each keypoint), which we note as KP-CLS. To go deeper, we conduct a study on the influence of such an alternative solution. As shown in Tab. 4, by comparing Row 2 with Row 1, we observe such a solution strengthens the quality perception of the model, resulting in performance improvement on score regression metrics. However, without explicit alignment with the keypoint-level coaching commentary, Bi-CLS falls short in understanding deeper relationships between the action execution and keypoints, resulting in a clear performance gap between our keypoint-level alignment, especially on commentary generation.
- Impacts of the Keypoint Query Augmentation. As discussed in Sec. 4.4, we construct keypoint queries by augmenting the keypoint text features (K.P.Text) with keypoint-type prompts (K.P.Type) and quality-related prompts (Quality). As shown in Tab. 5, we gradually ablate the components from the augmentation process, and the performance deterioration shows the contribution of each component.
- Impacts of videos from various views. As discussed in Sec. 5.2, we train TechCoach with both Egocentric and Exocentric videos. Here we conduct an experiment on the influence of the training videos from different views. As shown in Tab. 6, Ego-Only and Exo-Only variants showcase comparable performance. By combining videos from different views, TechCoach achieves stronger performance, which shows that egocentric and exocentric viewpoints are complementary for action coaching.
5.5 Qualitative Results
To further illustrate the effectiveness of the proposed TechCoach, we provide qualitative results on the generated coaching commentary. As shown in Fig. 5, TechCoach is able to understand various technical keypoints of different actions and provide detailed commentary on what is done well (e.g., “the extension of the shooting arm and wrist action were also commendable”) and what can be improved (e.g., “improvements are needed in core engagement to prevent being pulled from the wall”) from the given action videos. More qualitative results are shown in the Appendix.
6 Conclusion
In this paper, we investigate Descriptive Action Coaching, a novel task that focuses on providing coaching feedback on what is done well and what can be improved from an action execution. To achieve this, we propose TechCoach, a new framework that incorporates keypoint-level reasoning into the action coaching process, just as human coaches do. To support this task, we devise an automated pipeline for constructing the EE4D-DAC dataset which features clean and detailed coaching commentary on both keypoint and instance levels. Extensive experiments highlight the strong performance of the proposed TechCoach and demonstrate the effectiveness and the necessity of the proposed keypoint-level reasoning. We expect the proposed new task, dataset, and method provide a new perspective for advancing current AQA to a more explainable and practical scenario.
Appendix A Appendix
In this appendix, we provide more details about the evaluation metrics, implementations, and qualitative results in Appendix B. Subsequently, more details about the EE4D-DAC dataset are introduced in Appendix C, including examples of general technical keypoints, hierarchical coaching commentary, prompts used for LLM-based annotation collection, and other details. Finally, we discuss the limitations and potential directions for future work in Appendix D.
Appendix B More Details about Experiments
B.1 More Details about the Evaluation Metrics
Spearman’s rank correlation coefficient ($\rho$) [38] is the most widely used metric to evaluate models in the field of AQA. It measures the correlation between the predicted score series and the ground-truth score series, which is defined as follows:
$\rho=\frac{\sum^{N}_{i=1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum^{N}_{i=1}(x_{i}-\bar{x})^{2}\sum^{N}_{i=1}(y_{i}-\bar{y})^{2}}},$ (4)
where $x$ and $y$ are the rankings of the two series.
Relative-L2 Distance (RL2) [50] evaluates the numerical differences between the predicted and ground-truth scores, which can be written as:
$RL2=\frac{1}{N}\sum^{N}_{n=1}(\frac{|s_{n}-\hat{s}_{n}|}{s_{max}-s_{min}})^{2},$ (5)
where $s_{n}$ and $\hat{s}_{n}$ indicate the ground-truth and predicted scores, respectively. $s_{max}$ and $s_{min}$ indicate the highest and lowest scores for an action instance.
BLEU [25] quantifies the overlap between the generated sentences and the reference sentences by comparing n-grams (contiguous word sequences) shared between them. In this work, we compute the 1-gram-based BLEU score to evaluate the models.
METEOR [4] improves upon BLEU by introducing flexible word matching (e.g., synonyms and stemming) to better capture meaning beyond exact matches. Additionally, it adopts a recall-oriented approach to ensure all reference words are considered, and incorporates word order penalties to evaluate fluency and coherence.
CIDEr [37] is originally proposed for image captioning task. It measures the similarity between generated and reference texts using n-gram overlap weighted by TF-IDF. This ensures important but less common phrases are emphasized, rewarding consensus among references while penalizing overly generic outputs.
BERTScore [56] evaluates natural language generation by comparing the semantic similarity between generated and reference texts using contextual embeddings from pre-trained BERT-like models. Unlike traditional metrics that rely on n-gram overlap, BERTScore computes token-level similarity based on meaning, capturing nuances like synonyms and paraphrasing. In our work, we use the DeBERTa-xlarge [11] to extract token-level embeddings as it shows great correlation with human evaluation.
Figure 6: The detailed prompts for LLM-based evaluation. Zoom in for best view.
Figure 7: An overview of the Spatial-Temporal Context Enhancer. Here the $\oplus$ indicates the element-level adding operation.
GPT-M and GPT-Q are the LLM-based metrics proposed to evaluate whether the generated commentary shares the same technical details and the same coaching opinions as the ground-truth commentary. We employ the GPT-4o-mini to determine the number of both-mentioned technical details (i.e., keypoints) and the shared coaching opinions. Detailed prompts are provided in Fig. 6. After that, the GPT-M score is calculated as the averaged proportion of the both-mentioned technical details to those mentioned by ground truth across all testing instances. The GPT-Q score is calculated as the averaged proportion of the technical details sharing the same coaching opinions to the both-mentioned technical details.
B.2 More Implementation Details
Spatial-Temporal Context Enhancer. As discussed in Sec. 4.3 of the main paper, inspired by TimeSformer [5], we adopt a Spatial-Temporal Context Enhancer to enhance the spatial-temporal context across the clip-level features. An overview of this module is shown in Fig. 7. Specifically, we use a linear projector to map the clip-level features $F_{clip}$ to the hidden dimension $D$, and feed them into $N_{st}$ Transformer blocks that include spatial- and temporal-attention layers along with the FFN layers. In our implementation, the $N_{st}$ is set to 4.
Other details of TechCaoch. The visual backbone and text encoder are initialized by the pretrained weights and frozen during training. For the visual backbone, we use the pretrained weights of InternVideo2-stage3 [41]. Following [17, 54], we employ AdamW [18] as the optimizer, and utilize a learning rate warm-up during the early 10% training steps followed by linear decay. $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 and 0.3, respectively. Due to the characteristics of EgoExo4D [10] and the annotation pipeline of EE4D-DAC, it is not possible to ensure each instance includes rounded commentary on every possible technical keypoints (e.g., in some cases the coaching commentary would not mention the execution about the head & eyes). Therefore, we make some adjustments on the prompts or instructions for the compared models to guide the models to focus on those technical keypoints mentioned in the ground-truth Commentary. For TechCoach, we employ zero-masking on the keypoint-related quality embeddings corresponded to a specific keypoint only if both the strength and weakness on this keypoint are unmentioned in the ground truth. Note that such a design does NOT leak the action quality information, and similar adjustments are applied on the compared models [54, 16, 14, 41, 6] for fair comparisons.
Details about USDL [35]. USDL proposes to convert the score regression problem to a distribution prediction problem. We primarily follow the official implementation. We make adjustments on the output dimension to align with the score range in our experiments.
Details about CoRe [50] and TPT [3]. These two methods adopt the Contrastive Regression paradigm to generate action quality scores by predicting the score differences between the query instance and multiple exemplar instances. Our experiments are conducted based on the official implementations of these methods. The exemplars are randomly selected from the training instances with the same action task (e.g., Reverse Layup) as the query instance. The number of selected exemplars is set to 10.
Details about SwinBERT [17]. SwinBERT proposes a general framework for video captioning. In addition to the modification of the visual backbone, we also modified the length of masking tokens during training to ensure consistency with our TechCoach. Additionally, following [54], we add an extra full-connected layer to regress the score.
Details about PGMI [54]. PGMI is a prompt-guided framework for Multi-Task Learning AQA paradigm. Our experiments on PGMI are conducted using the official implementation. We adjust the number of score groups to 10 as the score range in our experiments is in [0, 10]. Besides, we modify the input templates for the text generator to include the score information and the technical dimensions that should be focused on. Specifically, the template is adjusted as follows:
The player is practicing <Action_Task>, and gets a score of <Predicted_Score>. Pay attention to the execution of <Technical_Dimensions>”,
where <Action_Task> is the action task, <Predicted_Score> indicates the predicted score by the model, and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground truth.
Details about the Compared MLLMs [16, 14, 41, 6]. We follow the official guidelines to perform zero-shot evaluation on the compared MLLMs. A general language instruction for these methods is as follows:
Suppose you are an <Action_Task>coach. Watch the given video, and answer a paragraph of commentary about what is done well in the execution and what can be improved. Your answer should be like: <Answer_Example>. Pay attention to the execution of <Technical_Dimensions>”,
where <Action_Task> is the action task (e.g., Reverse Layup), <Answer_Example> is an example of the answer (selected from training set), and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground-truth commentary.
B.3 More Qualitative Results
As a supplement to Fig. 5 in the main paper, we present additional qualitative results on the generated coaching commentary in Fig. 8, further illustrating the effectiveness of our TechCoach. Moreover, Fig. 9 demonstrates that our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details on various action executions.
Appendix C More Details about EE4D-DAC Dataset
In this section, we show the examples of the obtained general technical keypoints, hierarchical coaching commentary, prompts for annotation collection, and other details.
General Technical Keypoints. Completing a skilled physical action typically requires coordination across multiple body parts and objects. Thus, we define technical keypoints as language instructions spanning seven technical dimensions: 1) Head & Eyes; 2) Torso & Core; 3) Arms & Elbows; 4) Wrists & Hands; 5) Legs & Knees; 6) Feet & Heels; and 7) Human-Object Interaction. Tab. 7, Tab. 8 and Tab. 9 provide examples of the general technical keypoints for specific action tasks.
Hierarchical Coaching Commentary. Our EE4D-DAC dataset features hierarchical coaching commentary on both keypoint level and instance level. Fig. 10 and Fig. 11 show examples of the hierarchical coaching commentary.
Detailed Prompt for Annotation Collection. As described in Sec. 3.2 of the main paper, we use LLM (GPT-4o) to generate general technical keypoints and hierarchical coaching commentary. Additionally, as mentioned in Sec. 3.3, we prompt the LLM (GPT-4o) to compute the Coach-Score for commentary provided by various datasets. The detailed prompts are illustrated in Fig. 12.
Manual Check on Annotations. After obtaining the annotations with LLM, we manually check the evaluation set to ensure the annotation quality.
Appendix D Limitations and Future Work
Although experiments showcase the effectiveness of the proposed TechCoach, some limitations remain and warrant further exploration.
Hierarchical Technical Keypoints. In this work, for a given action task, we assume that each body part (or the dimension of human-object interaction) is associated with one paragraph of technical keypoints. While this assumption provides a general framework for the initial exploration of technical keypoint-aware DAC, technical keypoints could be further constructed hierarchically. For instance, when performing Mid-range Shooting, the shooting hand and guide hand play distinct roles in determining shooting accuracy and stability. Separating the technical keypoint for high-level body parts (e.g., wrists & hands) into finer sub-parts (e.g., shooting hand and guide hand) could further guide the model to focus on fine-grained and specific aspects in an action execution.
Pose-Assisted DAC. In this work, we focus on video-based DAC. While the Ego-Exo input enriches the visual context, there are cases where the egocentric video lacks sufficient visual information, and the human subject in the exocentric video appears relatively small. A valuable auxiliary modality to address this limitation is body pose. Since the original EgoExo4D [10] dataset provides frame-aligned 3D body pose data, incorporating pose sequences into the DAC process represents a promising direction for future research.
Figure 8: More qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors.
Figure 9: More qualitative results on the generated coaching commentary. Our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details (e.g., bending the knee) across various action executions. Correctly matched parts are highlighted in colors, and the corresponding regions are marked with dashed red circle. Some frames are zoomed in for better visibility.
Technical Dimensions General Technical Keypoints
Head & Eyes The player should keep the head up and eyes focused on the basket to maintain good court awareness and accuracy.
Torso & Core The player should engage the core to maintain balance and stability throughout the movement, allowing for smooth and controlled layups.
Arms & Elbows The player should use the arms to alternate the ball from one hand to the other while keeping the elbows close to the body to ensure control and precision during the layup.
Wrists & Hands The player should use the wrists to flick the ball off the backboard and into the basket, ensuring a soft touch with the hands for better control.
Legs & Knees The player should bend the knees to generate power for jumping, using the legs to propel themselves upward for the layup.
Feet & Heels The player should position the feet shoulder-width apart for balance, pushing off the heels to jump and land softly to maintain control.
Human-Object Interaction The player should practice alternating layups from one side of the basket to the other, using the backboard as a guide to ensure the ball goes into the basket consistently.
Table 7: General Technical Keypoints for Basketball Drills - Mikan Layup.
Technical Dimensions General Technical Keypoints
Head & Eyes The climber should keep the head up and eyes focused on the next hold. Always plan your next few moves in advance and be aware of your surroundings to maintain good spatial awareness.
Torso & Core The climber should engage the core to maintain balance and stability. A strong core helps in keeping the body close to the wall, reducing the effort needed from the arms.
Arms & Elbows The climber should keep the arms slightly bent rather than fully extended to prevent fatigue. Use the arms to pull the body upwards and keep movements controlled and deliberate.
Wrists & Hands The climber should use an open hand grip when possible, conserving energy. Make sure to grip holds firmly and use your fingers to feel for secure positions on the rock.
Legs & Knees The climber should use the legs to push the body up the wall, which helps conserve energy in the arms. Keep the knees slightly bent to absorb movements and maintain flexibility.
Feet & Heels The climber should place the feet carefully on holds, using the toes for precision. Heels can be used to hook onto holds for added stability when necessary.
Human-Object Interaction The climber should always check the stability of holds before putting the full weight on them. Use both hands and feet to distribute weight evenly and move smoothly from one hold to the next.
Table 8: General Technical Keypoints for Rock Climbing.
Technical Dimensions General Technical Keypoints
Head & Eyes The player should keep the head up and eyes focused on the ball at all times. This helps in tracking the ball’s movement and anticipating its next position.
Torso & Core The player should engage the core muscles to maintain balance and stability. A strong core helps in controlling the body’s movements and maintaining proper posture.
Arms & Elbows The player should keep the arms slightly bent and relaxed at the sides to help maintain balance. The arms can be used to make minor adjustments to the body’s position.
Wrists & Hands The player should use the hands to help with balance but should avoid using them to touch the ball. The wrists should be relaxed to allow natural movement.
Legs & Knees The player should keep the knees slightly bent to absorb the impact of the ball and to help with quick adjustments in positioning. This flexibility is crucial for maintaining control.
Feet & Heels The player should use the top of the feet to juggle the ball, aiming to keep it in the air with gentle, controlled touches. The heels should be lifted slightly off the ground to allow for quick movements.
Human-Object Interaction The player should focus on making consistent, light touches with the ball to keep it in the air. Each touch should be calculated to keep the ball at a manageable height and within close proximity to the body.
Table 9: General Technical Keypoints for Soccer Drills - Juggling.
Figure 10: An example for hierarchical coaching commentary. Zoom in for best view.
Figure 11: An example for hierarchical coaching commentary. Zoom in for best view.
Figure 12: The detailed prompts for generating General Technical Keypoints (Upper), Hierarchical Coaching Commentary (Middle), and Coach-Scores (Lower). “<Action_Domain>”: The action domain of the instance (e.g., Basketball, Soccer, Roch Climbing). “<Action_Domain>”: The action task of the instance (e.g., Reverse Layup). “<General_Technical_Keypoints>”: The generated general technical keypoints for a specific action task. “<Original_Expert_Commentary>”: The original expert commentary on the instance given by EgoExo4D [10]. Zoom in for best view.

### v2

@@
-TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching
+TechCoach: Towards Technical-Point-Aware Descriptive Action Coaching
 Abstract
-To guide a learner to master the action skills, it is crucial for a coach to 1) reason through the learner’s action execution and technical keypoints, and 2) provide detailed, understandable feedback on what is done well and what can be improved.
-However, existing score-based action assessment methods are still far from this practical scenario.
-To bridge this gap, we investigate a new task termed Descriptive Action Coaching (DAC) which requires a model to provide detailed commentary on what is done well and what can be improved beyond a quality score from an action execution.
-To this end, we construct a new dataset named EE4D-DAC.
-With an LLM-based annotation pipeline, our dataset goes beyond the existing action assessment datasets by providing the hierarchical coaching commentary at both keypoint and instance levels.
-Furthermore, we propose TechCoach, a new framework that explicitly incorporates keypoint-level reasoning into the DAC process.
-The central to our method lies in the Context-aware Keypoint Reasoner, which enables TechCoach to learn keypoint-related quality representations by querying visual context under the supervision of keypoint-level coaching commentary.
-Prompted by the visual context and the keypoint-related quality representations, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score.
-Combining all of these, we build a new benchmark for DAC and evaluate the effectiveness of our method through extensive experiments.
-Data and code will be publicly available.
-footnotetext: *: Corresponding author. $\dagger$: Equal contributions.
+To guide a learner in mastering action skills, it is crucial for a coach to 1) reason through the learner’s action execution and technical points (TechPoints), and 2) provide detailed, comprehensible feedback on what is done well and what can be improved.
+However, existing score-based action assessment methods are still far from reaching this practical scenario.
+To bridge this gap, we investigate a new task termed Descriptive Action Coaching (DescCoach) which requires the model to provide detailed commentary on what is done well and what can be improved beyond a simple quality score for action execution.
+To this end, we first build a new dataset named EE4D-DescCoach.
+Through an automatic annotation pipeline, our dataset goes beyond the existing action assessment datasets by providing detailed TechPoint-level commentary.
+Furthermore, we propose TechCoach, a new framework that explicitly incorporates TechPoint-level reasoning into the DescCoach process.
+The central to our method lies in the Context-aware TechPoint Reasoner, which enables TechCoach to learn TechPoint-related quality representation by querying visual context under the supervision of TechPoint-level coaching commentary.
+By leveraging the visual context and the TechPoint-related quality representation, a unified TechPoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score.
+Combining all of these, we establish a new benchmark for DescCoach and evaluate the effectiveness of our method through extensive experiments.
+The data and code will be made publicly available.
+footnotetext: $\dagger$: Corresponding author. *: Equal contributions.
 1 Introduction
@@
 Upper: Given an action demonstration, a human coach will provide coaching comments on what is done well and what can be improved.
-Lower: To provide precise feedback, the coach incorporates Keypoint-level Reasoning into the action coaching process, i.e., keep the technical keypoints (e.g., The player should extend the arms fully as they bring the ball under the basket) in mind and determine the keypoint-level action quality (e.g., The athlete performs beautifully with the right arm extension) of the action.
+Lower: To provide precise feedback, the coach incorporates TechPoint-level Reasoning into the action coaching process, i.e., keep the technical points (e.g., The player should extend the arms fully as they bring the ball under the basket) in mind and determine the technical-point-level action quality (e.g., The athlete performs beautifully with the right arm extension) of the action.
 Such a reasoning process motivates us to propose our TechCoach to address Descriptive Action Coaching.
@@
 1).
-However, existing works pose AQA as a score-regression [28, 35, 50] or pairwise-ranking [9, 55, 12] problem, limiting the applicability of the models to more practical scenarios.
+However, existing works formulate AQA as a score-regression [31, 38, 54] or pairwise-ranking [9, 60, 13] problem, limiting the applicability of the models to more practical scenarios.
 - Reasoning process.
-To provide precise feedback, a coach will keep the technical keypoints (e.g., The player should extend the arms fully in Reverse Layup) in mind and determine the keypoint-level action quality from action execution (see the lower part of Fig.
+To provide precise feedback, a coach will keep the technical points (TechPoints) in mind (e.g., The player should extend the arms fully in Reverse Layup) and determine the TechPoint-level action quality from action execution (see the lower part of Fig.
 1).
-However, few existing works have incorporated such a keypoint-level reasoning process into AQA.
-While recent works [39, 15] propose to address this issue by recognizing the existence of keypoint-level mistakes, such a paradigm still limits deeper exploration of the connections between the action execution and the keypoint-level action quality (i.e., detailed strength/weakness on a technical keypoint).
-To bridge the gap between the field of AQA and real-world action coaching scenarios, we investigate a new task termed Descriptive Action Coaching (DAC), which requires the model to provide detailed commentary on what is done well and what can be improved, beyond merely assigning a quality score to an action execution.
-Such a task is challenging because to provide such a detailed commentary, the model not only needs to understand fine-grained action execution (e.g., the player looks down when performing Reverse Layup), but also build explicit connections between the execution with the action quality (e.g., it is a weakness to look down at the ball as it will miss the target).
-To facilitate the studies on DAC, we construct EE4D-DAC, a new dataset that not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint and instance levels.
-Specifically, we first source the action videos from the recently proposed EgoExo4D [10] dataset.
-Subsequently, an LLM-driven automatic annotation pipeline is proposed to progressively obtain the general technical keypoints and hierarchical coaching commentary.
-The construction of the EE4D-DAC dataset allows us to build Technical Keypoint-aware Descriptive Action Coach (TechCoach), a new DAC framework which explicitly incorporates keypoint-level reasoning into the action coaching process.
-The key idea of TechCoach is to imitate the real-world coaching scenario, where the coach first reasons the action quality on each technical keypoint, and then provides feedback by incorporating the keypoint-level reasoning results and visual context.
-To this end, TechCoach includes a Context-aware Keypoint Reasoner, which queries the visual context with general technical keypoints to obtain keypoint-related quality representations.
-To ensure such representations carry the action quality information, a keypoint-level alignment loss is proposed to align the representations with the keypoint-level coaching commentary.
-After that, a Unified Keypoint-aware Action Assessor (KA2) is utilized to regress the action quality score and generate overall coaching commentary.
-In this process, a progressive action coaching attention mask is adopted to guide KA2 to integrate information starting with the visual context level, followed by the keypoint level, and ultimately the decision (i.e., score-prediction and commentary-generation) level.
-Combining the EE4D-DAC dataset and TechCoach, we can construct a new benchmark to expand the field of AQA.
-Extensive experiments demonstrate that: 1) Our TechCoach achieves state-of-the-art performance among the compared methods, including task-specific models and general MLLMs; 2) The Context-aware Keypoint Reasoner is a core design and cannot be simply replaced by other solutions (e.g., replacing the Keypoint-level Alignment with a classification-based solution that classify whether the execution shows strength or weakness on each keypoint).
-In summary, the contributions of this work can be summarized as: 1) We construct a new dataset termed EE4D-DAC to address Descriptive Action Coaching.
-To our knowledge, EE4D-DAC is the first dataset that includes hierarchical coaching commentary on both keypoint level and instance level.
-2) We develop TechCoach, a new method to address DAC by explicitly collaborating keypoint-level reasoning into the action coaching process.
-3) We construct a new benchmark on DAC.
-Experiments not only evaluate the effectiveness of the proposed approach but also provide a starting point for future work in this field.
-Figure 2: An overview of the LLM-driven automatic annotation pipeline for the EE4D-DAC dataset.
-We first prompt the LLM to generate general technical keypoints for the given action task (Orange $\rightarrow$).
-Subsequently, we ask the LLM to summarize the keypoint-level and instance-level commentary according to the general technical keypoints and the original expert commentary (Blue $\rightarrow$).
-Information about various keypoints is highlighted in colors.
+However, few existing works have incorporated such a TechPoint-level reasoning process into AQA.
+While recent works [42, 16] propose to address this issue by recognizing the existence of TechPoint-level mistakes, such a paradigm still limits deeper exploration of the connections between the action execution and the TechPoint-level action quality (i.e., detailed strength/weakness on a TechPoint).
+To bridge the gap between the field of AQA and real-world action coaching scenarios, we investigate a new task termed Descriptive Action Coaching (DescCoach), which requires the model to provide detailed commentary on what is done well and what can be improved, beyond merely assigning a quality score to an action execution.
+This task is challenging because to provide such a detailed commentary, the model not only needs to understand fine-grained action execution (e.g., the player looks down when performing Reverse Layup), but also establish explicit connections between the execution with the action quality (e.g., it is a weakness to look down at the ball as it will miss the target).
+To facilitate the studies on DescCoach, we construct EE4D-DescCoach, a new dataset that not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both TechPoint and instance levels.
+Specifically, we first source action videos from the recently proposed EgoExo4D [10] dataset.
+Subsequently, we design an LLM-driven automatic annotation pipeline to progressively obtain the general TechPoints and hierarchical coaching commentary.
+To our knowledge, no existing AQA dataset provides detailed TechPoint-level coaching commentary.
+Compared to existing datasets with score-based or binary TechPoint-level annotations [16, 25], we believe the TechPoint-level commentary provides more clear and detailed clues for reasoning about action quality and providing instance-level coaching commentary.
+The construction of the EE4D-DescCoach dataset allows us to build TechPoint-Aware Descriptive Action Coach (TechCoach), a new DescCoach framework which explicitly incorporates TechPoint-level reasoning into the action coaching process under the supervision of TechPoint-level coaching commentary, and delivers feedback by incorporating the reasoning results and visual context.
+To this end, TechCoach includes a Context-aware TechPoint Reasoner, which queries the visual context with general TechPoints to obtain TechPoint-related quality representations.
+To ensure these representations carry the action quality information, a TechPoint-level alignment loss is proposed to align the representations with the TechPoint-level coaching commentary.
+After that, a Unified TechPoint-aware Action Assessor (TA2) is employed to regress the action quality score and generate overall coaching commentary.
+Throughout this process, a progressive action coaching attention mask guides TA2 to integrate information starting with the visual context level, followed by the TechPoint level, and ultimately the decision level (i.e., score-prediction and commentary-generation).
+Combining the EE4D-DescCoach dataset and TechCoach, we establish a new benchmark to expand the field of AQA.
+Extensive experiments demonstrate that: (1) Our TechCoach achieves state-of-the-art performance among the compared methods, including task-specific models and general MLLMs; (2) The Context-aware TechPoint Reasoner is a core design to enhance action quality perceiving ability and cannot be simply replaced by naive alternative solutions (e.g., replacing the TechPoint-level alignment by classifying whether the execution shows strength or weakness on each TechPoint).
+The contributions of this work can be summarized as: (1) We introduce EE4D-DescCoach, a new dataset specifically designed for Descriptive Action Coaching.
+The hierarchical coaching commentary, particularly at the TechPoint level, establishes EE4D-DescCoach as a unique resource among existing AQA datasets.
+(2) We develop TechCoach, a new method that effectively integrates TechPoint-level reasoning into the action coaching process through supervision from TechPoint-level coaching commentary.
+(3) We construct a new benchmark on DescCoach.
+Experiments not only validate the effectiveness of the proposed approach, but also lay the foundation for future research in this field.
+Figure 2: An overview of the LLM-driven automatic annotation pipeline for the EE4D-DescCoach dataset.
+We first prompt the LLM to generate general TechPoints for the given action task (Orange $\rightarrow$).
+Subsequently, we ask the LLM to summarize the TechPoint-level and instance-level commentary according to the general TechPoints and the original expert commentary (Blue $\rightarrow$).
+Information about various TechPoints is highlighted in colors.
 Zoom in for best view.
 2 Related Works
-Action Quality Assessment (AQA) [43, 35, 50, 24, 47, 45, 3, 44, 29, 51, 27, 54, 30, 15, 22, 21, 57] aims at evaluating how well an action is performed.
-Most existing works pose AQA as a score regression task, and approach this task in many ways, e.g., introducing new formulations [43, 35, 50], augmenting human-body representations [24, 47], exploring temporal structures[45, 3, 44], collaborating multi-modal information [29, 51, 42].
-Although impressive progress has been achieved, regressing a single score is still far from real-world action coaching scenarios.
-While [27, 54] propose to generate commentary in addition to the score, the commentary comes from the live comment, focusing on action and score rather than skill guiding and coaching.
-Recently, considering it is crucial for a coaching model to reason through various technical keypoints, [30, 15, 39] propose to detect mistakes on various body parts or technical keypoints.
-[22, 21] introduce a scoring rubric to inject judging rules into the score prediction process.
-However, This branch of methods is still limited by the binary or score-only feedback.
-Notably, [23] incorporates multiple models together with judging rules to provide formatted feedback for diving judging, but requires complex manual design on a single task.
+Action Quality Assessment (AQA) [41, 20, 53, 62] aims at evaluating how well an action is performed.
+Most existing works pose AQA as a score regression task, and approach this task in many ways, e.g., introducing new formulations [46, 38, 54], augmenting human-body representations [27, 50], modeling with insufficient or continual data [55, 17, 63], exploring temporal structures[48, 3, 47], collaborating multi-modal information [32, 56, 45].
+Although impressive progress has been achieved, regressing the final score is still far from real-world action coaching scenarios.
+MTL-AQA[30] and NAE[59] propose to generate commentary in addition to the score, but the commentary comes from live comment, focusing on action and score rather than skill guiding and coaching.
+NS-AQA[26] combines multiple models together with judging rules to provide formatted feedback for diving judging, but requires complex manual design on a single task.
 Differently, our work focuses on generating free-from commentary on what is done well and what can be improved, which is more aligned with practical coaching scenarios.
-To achieve this, we construct a new dataset that features detailed coaching commentary on both keypoint and instance levels.
-Furthermore, we propose a new general framework that collaborates key-point reasoning into coaching process.
-Video Captioning [34, 46, 17, 40, 19, 49] is a long-lasted task that requires a model to generate language caption with video frames as input.
+To this end, we construct a new dataset that features hierarchical (especially the TechPoint-level) detailed coaching commentary.
+Comparisons with existing datasets are shown in Sec.
+3.1.
+Furthermore, we propose a general framework that collaborates TechPoint-level reasoning into the coaching process.
+Different from existing works [25, 24, 33, 16, 42] that use binary or score-based supervised pipeline to introduce score rubric or detect body-part level mistakes, TechCoach represents the first approach to learn action quality-related representation under the supervision of TechPoint-level commentary.
+Experiments further validate the effectiveness of the unique design against other naive solutions.
+Video Captioning [37, 49, 19, 43, 22, 52] is a long-standing task that requires a model to generate a language caption with video frames as input.
 Early studies focus on designing stronger architectures [19, 43, 1, 36, 51] or utilizing video-language pretraining [37, 22, 52, 2, 35] to model relationships between video and language.
 Recently, Multimodal Large Language Models (MLLMs) [18, 44, 15, 23, 14, 7] show strong generalization ability across various video-language tasks such as video content description and video question answering.
-In this paper, we focus on Descriptive Action Coaching, a rarely explored but challenging scenario that requires video-language models to collaborate fine-grained video content and domain knowledge to provide detailed coaching commentary.
-We provide a new testbed for today’s video-language models and show that they still struggle to handle such a practical scenario.
-Besides, we propose a new keypoint-guided coaching model and evaluate its effectiveness through extensive experiments.
-3 EE4D-DAC Dataset
-To enable studies on Descriptive Action Coaching, we construct a new dataset termed EE4D-DAC.
-Different from existing action assessment datasets, EE4D-DAC not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint level and instance level.
+However, rare exploration has been conducted on whether these models could provide coaching commentary about what is done well and what can be improved.
+In this work, we provide a new testbed for today’s video-language models and reveal that they still struggle to address such a practical scenario.
+Besides, we propose a new TechPoint-guided coaching framework and evaluate its effectiveness through extensive experiments.
+3 EE4D-DescCoach Dataset
+To enable studies on Descriptive Action Coaching, we establish a new dataset termed EE4D-DescCoach.
+EE4D-DescCoach not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both technical-point (TechPoint) level and instance level.
 We will introduce the data source in Sec.
@@
 3.2, and provide comparisons with existing datasets in Sec.
-3.3 respectively.
+3.3.
 3.1 Data Source
 We source data from the recently proposed EgoExo4D [10] dataset.
-Except for various ego-exo action videos, EgoExo4D also contains execution rating (from 1 to 10) and time-aligned free-form spoken commentary (e.g.
+In addition to ego-exo action videos, EgoExo4D also contains execution rating (1 to 10) and time-aligned free-form spoken commentary (e.g.
 “Yeah, see how she’s having to look down.
@@
 By doing so, we obtain 4843 unique video instances, spanning about 10.8 hours.
-Datasets Years Data Unique Unique KeyPoint Level Judgment Instance Level Judgment
+{NiceTabular}
+l|c|ccc|ccc|ccc Datasets Public. Data Unique Unique TechPoint Level Judgment   Instance Level Judgment
 Sources Instances Hours Exists Type Num Score Commentary Coach-Score
 Datasets w/o Language Annotations
-MIT-Skate [31] ECCV14 Olympics 150 7.3h $\times$ - - $\times$ -
-AQA-7 [26] WACV19 Olympics 1189 1.62h $\times$ - - $\times$ -
-Rhythmic Gymnastics [52] ACMMM20 Olympics 1000 23.4h $\times$ - - $\times$ -
-FineDiving [45] CVPR22 Olympics 3000 3.5h $\times$ - - $\times$ -
-LOGO [53] CVPR23 Olympics 200 11.4h $\times$ - - $\times$ -
-Skate-IRIS [22] ACMIUI23 MIT-Skate [31] 150 7.3h Score 1050 $\times$ -
-CPR-Coach [39] CVPR24 Live Saving 1416 7.7h Binary 18.4k $\times$ $\times$ -
+MIT-Skate [34] ECCV14 Olympics 150 7.3h $\times$ - - ✓ $\times$ -
+AQA-7 [29] WACV19 Olympics 1189 1.62h $\times$ - - ✓ $\times$ -
+Rhythmic Gymnastics [57] ACMMM20 Olympics 1000 23.4h $\times$ - - ✓ $\times$ -
+FineDiving [48] CVPR22 Olympics 3000 3.5h $\times$ - - ✓ $\times$ -
+LOGO [58] CVPR23 Olympics 200 11.4h $\times$ - - ✓ $\times$ -
+Skate-IRIS [25] ACMIUI23 MIT-Skate [34] 150 7.3h ✓ Score 1050 ✓ $\times$ -
+CPR-Coach [42] CVPR24 Live Saving 1416 7.7h ✓ Binary 18.4k $\times$ $\times$ -
 Datasets w/ Language Annotations
-MTL-AQA [27] CVPR19 Olympics 1412 1.7h $\times$ - - 1.94
-MTL-NAE [54] CVPR24 MTL-AQA [27] 1412 1.7h $\times$ - - 2.94
-EgoExo-Fitness [15] ECCV24 Daily Fitness 913 4.6h Binary 7.8k 2.37
-EgoExo4D (Sports) [10] CVPR24 Daily Sports 1219 16.4h $\times$ - - 3.44
-EE4D-DAC(Ours) EgoExo4D [10] 4843 10.8h Commentary 25.1k 4.33
+MTL-AQA [30] CVPR19 Olympics 1412 1.7h $\times$ - - ✓ ✓ 1.94
+MTL-NAE [59] CVPR24 MTL-AQA [30] 1412 1.7h $\times$ - - ✓ ✓ 2.94
+EgoExo-Fitness [16] ECCV24 Daily Fitness 913 4.6h ✓ Binary 7.8k ✓ ✓ 2.37
+EgoExo4D (Sports) [10] CVPR24 Daily Sports 1219 16.4h $\times$ - - ✓ ✓ 3.44
+EE4D-DescCoach (Ours) EgoExo4D [10] 4843 10.8h Commentary 25.1k 4.33
 Table 1: Comparison to popular Action Assessment datasets.
-Our EE4D-DAC dataset is the first dataset that contains detailed coaching commentary on both key-point level and instance level.
+Our EE4D-DescCoach dataset is the first dataset that contains hierarchical (especially TechCoach-level) detailed coaching commentary.
 Coach-Score: A metric that evaluates whether the provided commentary is suited for real-world coaching scenarios (See Sec.
@@
 3.2 Annotation Pipeline
-By going deeper into the expert commentary in EgoExo4D, it can be observed that: 1) The original expert commentary is highly colloquial and noisy, awaiting summarization before being used to build a coaching model.
-2) The expert commentary is highly associated with the general technical keypoints, making it possible to mine the relationships between the general technical keypoints and final coaching commentary.
-Based on the observations, we propose an LLM-based automatic annotation pipeline to obtain general technical keypoints, and hierarchical coaching commentary on both keypoint level and instance level, which is shown in Fig.
+By diving deeper into the expert commentary in EgoExo4D, it can be observed that: 1) The original expert commentary is highly colloquial and noisy, awaiting summarization before being used to build a coaching model.
+2) The expert commentary is highly associated with the general TechPoints, making it possible to mine the relationships between the general TechPoints and final coaching commentary.
+Based on the observations, we propose an LLM-based automatic annotation pipeline to obtain general TechPoints, and hierarchical coaching commentary on both TechPoint level and instance level, which is shown in Fig.
 2.
 We will bring more details as follows.
-- Collection of the General Technical Keypoints.
-In the general coaching-learning scenario, a coach will provide several technical keypoints in advance so that the skill learner can learn to follow.
-Considering that completing a skilled physical action always requires cooperation across multiple body parts and objects, we define a technical keypoint as a language instruction on one of the seven dimensions, including six body parts (e.g., head & eyes, arms & hands) and the human-objects interaction.
-To obtain technical keypoints for each skilled action task, we treat the LLM (GPT-4O) as a general coach.
+- General TechPoints Collection.
+In the general coaching-learning scenario, a coach will provide several TechPoints in advance so that the skill learner can learn to follow.
+Considering that completing a skilled physical action always requires cooperation across multiple body parts and objects, we define a TechPoint as a language instruction on one of the seven dimensions, including six body parts (e.g., head & eyes, arms & hands) and the human-objects interaction.
+To obtain TechPoints for each skilled action task, we treat the LLM (GPT-4o) as a general coach.
 As shown in the lower left part of Fig.
-2, by providing the task name (e.g., Basketball Drills - Reverse Layup) and detailed instructions to the LLM, we are able to obtain the general technical keypoints on the previously mentioned seven dimensions (e.g., For the dimension of “head & eyes”, we have the keypoint: “The player should keep the head up and eyes focused on the backboard and rim to accurately judge the distance and angle for the reverse layup.”).
-- Collection of Hierarchical Coaching Commentary.
-Subsequently, we further ask the LLM to mine the relationships between the technical keypoints and origin expert commentary and provide summarized hierarchical (i.e., keypoint level and instance level) coaching commentary about what is done well and what can be improved in the action execution.
+2, by providing the task name (e.g., Basketball Drills - Reverse Layup) and detailed instructions to the LLM, we are able to obtain the general TechPoints on the previously mentioned seven dimensions (e.g., For the dimension of “head & eyes”, we have the TechPoint: “The player should keep the head up and eyes focused on the backboard and rim to accurately judge the distance and angle for the reverse layup.”).
+- Hierarchical Coaching Commentary Collection.
+Subsequently, we further ask the LLM to mine the relationships between the TechPoints and origin expert commentary and provide summarized hierarchical (i.e., TechPoint level and instance level) coaching commentary about what is done well and what can be improved in the action execution.
 As shown in the right part of Fig.
-2, in this phase, LLM is asked to finish two tasks: (1) Review the general technical keypoints and the original expert commentary, and then summarize the strength and weakness in the execution corresponding to each keypoint.
+2, in this phase, LLM is asked to finish two tasks: (1) Review the general TechPoints and the original expert commentary, and then summarize the strength and weakness in the execution corresponding to each TechPoint.
 (2) Review the original expert commentary and the generated strengths and weaknesses, and then provide the overall commentary on the execution.
-To sum up, with the annotation pipeline, for each unique instance, we construct rich language annotations including the General Technical Keypoints, Keypoint-level Commentary, and Instance-level Commentary.
+To sum up, with the annotation pipeline, for each unique instance, we construct rich language annotations including the General TechPoints, TechPoint-level Commentary, and Instance-level Commentary.
 Moreover, for each instance we also obtain an average rating by averaging all the ratings provided by different experts.
 - Discussions.
-Note that in some cases, an execution could reveal both the strength and weakness aspects on one keypoint, and some may just reveal one of them.
+Note that in some cases, an execution could reveal both the strength and weakness aspects on one TechPoint, and some may just reveal one of them.
 Such a characteristic is inherited from EgoExo4D [10].
 3.3 Comparison with existing datasets
-We compare our dataset with related AQA datasets.
-As shown in Tab.
-1, the detailed keypoint-level commentary allows us to separate our EE4D-DAC dataset from the other datasets [22, 39, 15] with keypoint-level annotations on both type and scale.
+We compare our dataset with related AQA-related datasets.
+As shown in Sec.
+3.1, the detailed TechPoint-level commentary allows us to distinguish our EE4D-DescCoach dataset from the other related datasets [25, 42, 16].
 Furthermore, we conduct an evaluation to show the guiding ability of the instance-level commentary given by different datasets.
 Specifically, we prompt the ChatGPT to rate the commentary from 0 to 5 as the Coach-Score.
-As shown in the last column in Tab.
-1, commentary MTL-AQA [27] and MTL-NAE [54] is not suit for coaching scenario as the original text mainly focuses on the action and score rather than action guiding and skill improvement.
-For more details about the prompts, the pre-defined body parts, the examples and dataset comparison strategies, please refer to the Appendix.
+As shown in the last column in Sec.
+3.1, commentary MTL-AQA [30] and MTL-NAE [59] is not suit for coaching scenario as the original text mainly focuses on the action and score rather than action guiding and skill improvement.
+For more details about the prompts, the pre-defined body parts, the examples and quality ensurance, please refer to the Appendix.
 4 TechCoach
@@
 Given an action video, TechCoach first extracts visual context embeddings with a pre-trained visual backbone and a Spatial-Temporal Context Enhancer.
-After that, a Context-aware Keypoint Reasoner is adopted to learn keypoint-related quality embeddings by querying visual context under the supervision of keypoint-level coaching commentary.
-Finally, prompted by the visual context and the keypoint-related quality embeddings, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score.
-K.P.: Keypoints.
+After that, a Context-aware Tech-point Reasoner is adopted to learn Tech-point-related quality embeddings by querying visual context under the supervision of TechPoint-level coaching commentary.
+Finally, prompted by the visual context and the TechPoint-related quality embeddings, a unified TechPoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score.
+T.P.: TechPoints.
 Best viewed in color.
-Figure 4: An overview of the Context-aware Keypoint Reasoner. Best viewed in color.
+Figure 4: An overview of the Context-aware TechPoint Reasoner. Best viewed in color.
 4.2 Overview
-An overview of the proposed Technical Keypoint-aware Descriptive Action Coach (TechCoach) is shown in Fig.
+An overview of the proposed Technical-Point-aware Descriptive Action Coach (TechCoach) is shown in Fig.
 3.
-The key idea of TechCoach is to mimic the reasoning process of human coaches by incorporating keypoint-level reasoning into the action coaching pipeline.
+The key idea of TechCoach is to mimic the reasoning process of human coaches by incorporating TechPoint-level reasoning into the action coaching pipeline.
 Our framework consists of three parts.
 Firstly, given an input video $v$, a Visual Encoding Module will extract the visual context embedding.
-After that, a Context-aware Key-point Reasoner is adopted to take the visual context embedding and general technical keypoints as input to perceive the action quality on each keypoints and provide the key-point quality embeddings.
-Finally, a Unified Keypoint-aware Action Assessor is utilized to collaborate the visual context and keypoint-aware quality embeddings, then predict the final quality score and generate coaching commentary.
+After that, a Context-aware TechPoint Reasoner is adopted to take the visual context embedding and general technical points as input to perceive the action quality on each TechPoints and provide the TechPoint quality embeddings.
+Finally, a Unified TechPoint-aware Action Assessor is utilized to collaborate the visual context and TechPoint-aware quality embeddings, then predict the final quality score and generate coaching commentary.
 Discussions.
-Note that our model is based on the assumption that the general technical keypoints are available.
-We want to argue that this assumption is reasonable because: 1) It is aligned with real-world coaching scenarios that a coach is well-informed about the keypoints and able to reason through the keypoints and learner’s executions.
-2) Obtaining keypoints is not resource-intensive as we can prompt LLMs (e.g., ChatGPT) to provide the keypoints.
+Note that our model is based on the assumption that the general TechPoints are available.
+We want to argue that this assumption is reasonable because: 1) It is aligned with real-world coaching scenarios that a coach is well-informed about the TechPoints and able to reason through the TechPoints and learner’s executions.
+2) Obtaining TechPoints is not resource-intensive as we can prompt LLMs (e.g., ChatGPT) to provide the TechPoints.
 4.3 Visual Encoding
@@
 The details about the Spatial-Temporal Context Enhancer are presented in Appendix.
-4.4 Context-aware Key-point Reasoner
-To generate comprehensive coaching commentary for an action execution, an intuitive approach is to directly input the video (and the keypoint) features into the text generator.
-However, without explicit modeling, such an approach is unable to effectively perceive the relationship between visual context and the keypoints.
-To address this, TechCoach adopts a Context-aware Key-point Reasoner to learn keypoint-related quality embeddings by querying the video context embeddings under the supervision of the coaching commentary for each technical keypoint.
-An overview of Context-aware Keypoint Reasoner is shown in Fig.
+4.4 Context-aware TechPoint Reasoner
+To generate comprehensive coaching commentary for an action execution, an intuitive approach is to directly input the video (and the TechPoint) features into the text generator.
+However, without explicit modeling, such an approach is unable to effectively perceive the relationship between visual context and the TechPoints.
+To address this, TechCoach adopts a Context-aware TechPoint Reasoner to learn TechPoint-related quality embeddings by querying the video context embeddings under the supervision of the coaching commentary for each TechPoint.
+An overview of Context-aware TechPoint Reasoner is shown in Fig.
 4.
-- Keypoint Queries Construction.
-Without loss of generality, given the N general technical keypoints corresponding to the input video, we first extract the keypoint features $f_{kp}\in R^{N\times D}$ with a pre-trained text encoder (e.g., DeBERTa [11]) and a linear mapper.
-After that, we augment $f_{kp}$ by introducing two types of learnable prompts.
-Specifically, we use a group of keypoint-type prompts $f_{kt}\in R^{N\times D}$ to identify which dimension (e.g., head & eyes, human-object interaction) each keypoint feature belongs to.
+- TechPoint Queries Construction.
+Without loss of generality, given the N general TechPoints corresponding to the input video, we first extract the TechPoint features $f_{tp}\in R^{N\times D}$ with a pre-trained text encoder (e.g., DeBERTa [11]) and a linear mapper.
+After that, we augment $f_{tp}$ by introducing two types of learnable prompts.
+Specifically, we use a group of TechPoint-type prompts $f_{tt}\in R^{N\times D}$ to identify which dimension (e.g., head & eyes, human-object interaction) each TechPoint feature belongs to.
 Besides, we adopt two quality-related prompts $f_{q}\in R^{2\times D}$ to indicate the aspects of strength and weakness.
-Based on the definitions of the learnable prompts, the augmented keypoint features $f^{*}_{kp}\in R^{N\times 2\times D}$ are computed as:
-$f^{*}_{kp}[i,j]=f_{kp}[i]+f_{kt}[i]+f_{q}[j].\vspace{-0.2cm}$ (1)
-Such an augmentation allows us to inject the keypoint and quality information into the keypoint queries.
-- Keypoint-driven Quality Reasoning.
-Our goal here is to decouple the visual context features corresponding to the keypoints and quality aspects.
-To this end, we regard the augmented keypoint features $f^{*}_{kp}$ as the queries and visual context features $f_{v}$ as the keys and values, then utilize layers of multi-head cross attention module [36] to obtain the keypoint-related quality embeddings $f_{kq}\in R^{N\times 2\times D}$.
-$f_{kq}[i,j]$ indicates the action quality on keypoint $i$ and aspect $j$ ($1$ for the strength and $2$ for the weakness).
-To ensure $f_{kq}$ carries the quality information related to the action execution on the technical keypoints, our idea is to align the keypoint-related quality embeddings $f_{kq}$ with the features of the corresponding keypoint-level coaching commentary.
-To this end, a keypoint-level alignment loss is adopted, which can be written as:
-$L_{Align}=\sum^{N}_{i=1}\sum^{2}_{j=1}w_{kc}[i,j]*D(\psi(f_{kq}[i,j]),f_{kc}[i,j]),\vspace{-0.2cm}$ (2)
-where $f_{kc}$ indicates the features of keypoint-level commentary extracted by the text encoder; $\psi$ is a linear mapper that projects the $f_{kq}$ back to the dimension of the text features; $w_{kc}[i,j]$ is a binary scalar to filter those cases where commentary on some aspects is not provided (see Discussions in Sec.
-3.2); $D(\cdot,\cdot)$ indicates a distance measurement between two features, and we use the L2 Distance by default.
-4.5 Unified Keypoint-aware Action Assessor
-After obtaining the visual context embeddings and the keypoint-related quality embeddings, we employ a Transformer-based Unified Keypoint-aware Action Assessor (KA2) to jointly predict the action quality score and generate the overall coaching commentary.
-Specifically, the input sequence to the KA2 module is constructed by concatenating the following components: (1) The generated text tokens $T_{1:i-1}$; (2) a learnable score-prediction token $f_{s}$; (3) the flattened $f_{kc}$ as keypoint tokens; (4) the flattened $f_{v}$ as video tokens.
+Based on the definitions of the learnable prompts, the augmented TechPoint features $f^{*}_{tp}\in R^{N\times 2\times D}$ are computed as:
+$f^{*}_{tp}[i,j]=f_{tp}[i]+f_{tt}[i]+f_{q}[j].\vspace{-0.2cm}$ (1)
+Such an augmentation allows us to inject the TechPoint and quality information into the TechPoint queries.
+- TechPoint-driven Quality Reasoning.
+Our goal here is to decouple the visual context features corresponding to the TechPoints and quality aspects.
+To this end, we regard the augmented TechPoint features $f^{*}_{tp}$ as the queries and visual context features $f_{v}$ as the keys and values, then utilize layers of multi-head cross attention module [39] to obtain the TechPoint-related quality embeddings $f_{tq}\in R^{N\times 2\times D}$.
+$f_{tq}[i,j]$ indicates the action quality on TechPoint $i$ and aspect $j$ ($1$ for the strength and $2$ for the weakness).
+To ensure $f_{tq}$ carries the quality information related to the action execution on the TechPoints, our idea is to align the TechPoint-related quality embeddings $f_{tq}$ with the features of the corresponding TechPoint-level coaching commentary.
+To this end, a TechPoint-level alignment loss is adopted, which can be written as:
+$L_{Align}=\sum^{N}_{i=1}\sum^{2}_{j=1}w_{tc}[i,j]*D(\psi(f_{tq}[i,j]),f_{tc}[i,j]),\vspace{-0.4cm}$ (2)
+where $f_{tc}$ indicates the features of TechPoint-level commentary extracted by the text encoder; $\psi$ is a linear mapper that projects the $f_{tq}$ back to the dimension of the text features; $w_{tc}[i,j]$ is a binary scalar to filter those cases where commentary on some aspects is not provided (see Discussions in Sec.
+3.2); $D(\cdot,\cdot)$ indicates a distance measurement between two features and we use the L2 Distance by default.
+4.5 Unified TechPoint-aware Action Assessor
+After obtaining the visual context embeddings and the TechPoint-related quality embeddings, we employ a Transformer-based Unified TechPoint-aware Action Assessor (TA2) to jointly predict the action quality score and generate the overall coaching commentary.
+Specifically, the input sequence to the TA2 module is constructed by concatenating the following components: (1) The generated text tokens $T_{1:i-1}$; (2) a learnable score-prediction token $f_{s}$; (3) the flattened $f_{tq}$ as TechPoint tokens; (4) the flattened $f_{v}$ as video tokens.
 For generating commentary, we follow [19, 59] by employing a Mask-Token Prediction supervision ($L_{MTM}$) during training and using the Next-Token Prediction paradigm during inference.
 For score regression, the score-prediction token from the output layer is passed to a fully connected regressor, optimized using a Mean Squared Error (MSE) loss ($L_{MSE}$).
-Moreover, to guide the KA2 module in progressively integrating information in a sequence from low-level to high-level (i.e., starting with visual context level, then keypoint level, and finally the decision level), a progressive action coaching attention mask is further adopted on the KA2, which is shown in the lower right part of Fig.
+Moreover, to guide the TA2 module in progressively integrating information in a sequence from lower to higher level (i.e., starting with visual context level, then TechPoint level, and finally the decision level), a progressive action coaching attention mask is further adopted on the TA2, which is shown in the lower right part of Fig.
 3.
-Specifically, each keypoint token independently attends to itself and the video tokens.
-Besides, the score-prediction token attends to itself and all the keypoint and video tokens.
-For commentary generation, KA2 integrates all information by attending to the generated text tokens, the score-prediction token, all the keypoint tokens, and the video tokens.
+Specifically, each TechPoint token independently attends to itself and the video tokens.
+Besides, the score-prediction token attends to itself and all the TechPoint tokens and video tokens.
+For commentary generation, TA2 integrates all information by attending to the generated text tokens, the score-prediction token, all the TechPoint tokens, and the video tokens.
 To sum up, we train the TechCoach with the following overall loss function:
-$L=L_{MTM}+\lambda_{1}L_{MSE}+\lambda_{2}L_{Align},\vspace{-0.2cm}$ (3)
+$L=L_{MTM}+\lambda_{1}L_{MSE}+\lambda_{2}L_{Align},\vspace{-0.3cm}$ (3)
 where $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters.
-Methods Years Score Regression
-$\rho\uparrow$ RL2$\downarrow$
+{NiceTabular}
+l|c@  |c@   c@    Methods Visual Score Regression
+Backbone $\rho\uparrow$ RL2$\downarrow$
 Models infer with single instance
-InternVideo2-MLP - 69.75 4.70
-USDL [35] CVPR21 70.81 4.51
-SwinBERT [17] CVPR22 71.37 4.59
-PGMI [54] CVPR24 70.93 4.38
-TechCoach(Ours) 72.46 4.35
+InternVideo2-MLP IV2 [44] 69.75 4.70
+USDL [38] IV2 [44] 70.81 4.51
+SwinBERT [19] IV2 [44] 71.37 4.59
+PGMI [59] IV2 [44] 70.93 4.38
+TechCoach(Ours) IV2 [44] 72.46 4.35
 Models infer with extra exemplars
-CoRe [50] w/ 10 Exem. ICCV21 73.17 4.35
-TPT[3] w/ 10 Exem. ECCV22 73.77 4.29
-(a) Comparison with existing methods on score regression.
-Methods Years Commentary Generation
+CoRe [54] w/ 10 Exem. IV2 [44] 73.17 4.35
+TPT[3] w/ 10 Exem. IV2 [44] 73.77 4.29
 […]
-General Multimodal Large Language Models
-VideoLLaVA-7B [16] ACL24 22.94 1.79 9.80 53.76 16.62 14.82
-VideoChat2-7B [14] CVPR24 23.07 3.20 11.94 60.79 17.82 25.99
-InternVideo2-Stage3-8B [41] ECCV24 27.52 2.98 12.83 63.80 37.85 28.86
-InternVL2-8B [6] 2024 12.51 0.10 16.81 62.30 47.76 34.01
-Task-Specific Models
-SwinBERT[17] CVPR22 36.11 11.70 16.01 66.03 46.56 34.57
-PGMI [54] CVPR24 36.78 14.00 16.24 66.42 49.02 37.04
-TechCoach(Ours) - 37.06 14.62 16.39 66.89 50.44 38.15
 (b) Comparison with existing methods on coaching commentary generation.
@@
 “w/ 10 Exem.”: We gray out models that infer with 10 extra exemplars for fair comparisons.
+“Txt.Gen.”: Text Generator.
+“IN2”: InternVideo2.
 5 Experiments
@@
 - Dataset.
-We conduct experiments on EE4D-DAC to provide a new benchmark and evaluate the performance of the proposed TechCoach.
-Following the official split of EgoExo4D [10], we separate our EE4D-DAC into training and evaluation set (including 3769 and 1074 instances, respectively).
+To build new benchmark and evaluate the effectiveness of our TechCoach, we separate EE4D-DescCoach into training and evaluation set (with 3769 and 1074 instances) following the official split of EgoExo4D [10].
 - Evaluation Metrics.
@@
 For natural language generation, we adopt NLP metrics, including BERT [61], BLEU(B)[28], METEOR(M)[4] and CIDEr(C) [40] scores.
-Moreover, for better comparison with open-source MLLMs, we design two more LLM-based Metrics: (1) a Mention Score(GPT-M) to evaluate whether the generated commentary mentions the same technical details as in the ground truth; (2) a Quality Score(GPT-Q) to evaluate whether the generated commentary shares the same praises or improvement opinions as in the ground truth on those both-mentioned technical details.
-For more details about the evaluation metrics, please refer to the Appendix.
+Moreover, for better comparison with open-source MLLMs, we design two more LLM-based Metrics: (1) a Mention Score(GPT-M) to evaluate whether the generated commentary mentions the same technical details as in the ground truth; (2) a Quality Score(GPT-Q) to evaluate whether the generated commentary shares the same praises or improvement opinions as in the ground truth on those both-mentioned technical details (i.e., action quality perception ability).
+Please refer to Appendix for more details.
 5.2 Implementation Details
@@
 obtaining video features with a size of $16(T)\times 8(H)\times 8(W)$.
-We use the pretrained DeBERTa-xlarge [11] as the text encoder.
 We use both the ego- and exo-centric (best-exo in EgoExo4D) videos as input.
 We concatenate the multi-view visual features on channel dimension and feed them into a linear projector before going through the Spatial-Temporal Context Enhancer.
-For more implementation details, please refer to the Appendix.
+We use similar multi-modal Transformer encoder as in [59, 19] to integrate inputs and generate text.
+More implementation details are provided in the Appendix.
 5.3 Main Results
-We compare the proposed methods with two branches of baselines: (1) Task-Specific Models: We select AQA- and VideoCaption-specific baselines including USDL [35], CoRe [50], TPT[3], SwinBERT[17] and PGMI [54].
-For fair comparisons, we evaluate all domain-specific models by using the same visual backbone (i.e., InternVideo2) to extract features and the same fusion strategy on multi-view features.
-(2) General Multi-modal Large Language Models: We also select popular open-source MLLMs as the baselines, including VideoLLaVA [16], VideoChat2 [14], InternVideo2-Stage3 [41], and InternVL2 [6].
-For more details about the baseline models, please refer to the Appendix.
-The comparison results are shown in Tab.
-2 and we have the following observations:
-Variants Score Commentary
+As shown in Sec.
+4.5, we compare the proposed methods with two branches of baselines: (1) Task-Specific Models: We select task-specific baselines including USDL [38], CoRe [54], TPT[3], SwinBERT[19] and PGMI [59].
+For fair comparisons, we evaluate all task-specific models by using the same visual backbone (i.e., InternVideo2) to extract features and the same fusion strategy on multi-view features.
+(2) General Multi-modal Large Language Models: We compare popular open-source MLLMs including VideoLLaVA [18], VideoChat2 [15], InternVideo2-S3 [44], and InternVL2 [6] under zero-shot evaluation settings.
+We provide more details about the baselines in the Appendix.
+{NiceTabular} ccc|cc|cccc Variants Score Commentary
 […]
-70.78 4.45 - - - -
-- - 36.71 13.70 16.39 66.77
-70.23 4.77 36.42 14.26 16.09 65.72
-72.46 4.35 37.06 14.62 16.39 66.89
-Table 3: Ablation studies on training losses. The results show that keypoint-level alignment is a core design to achieve a better trade-off between score regression and coaching commentary generation.
-Variants Score Commentary
+✓ 70.78 4.45 - - - -
+✓ - - 36.71 13.70 16.39 66.77
+✓ ✓ 70.23 4.7736.42 14.26 16.09 65.72
+✓ ✓ ✓ 72.46 4.35 37.06 14.62 16.39 66.89
+Table 3: Ablation studies on training losses. TechPoint-level alignment is a core design to achieve a better trade-off between score regression and coaching commentary generation.
+Refer to caption Figure 5: A further illustration of the impact of TechPoint-level alignment loss.
+After adding $L_{Align}$, a much stronger improvement is observed on action quality perception ability (evaluated by GPT-Q Score).
+Please refer to Sec.
+5.4 for more analysis.
+- Score Regression: Compared with Direct Regression-based (i.e., InternVideo2-MLP and USDL) and Multi-Task Learning-based (i.e., SwinBERT, PGMI) AQA methods, TechCoach achieves the best performance on score regression.
+Note that CoRe and TPT adopt a Contrastive Regression framework that utilizes 10 extra exemplars during inference, thereby enhancing their score regression capability.
 […]
-w/o KP-Align 70.23 4.77 36.42 14.26 16.09 65.72
-KP-CLS 72.13 4.45 36.78 13.69 16.19 64.69
-KP-Align 72.46 4.35 37.06 14.62 16.39 66.89
-Table 4: Ablation studies on the alternative solutions of keypoint-level alignment.
-The results show that the proposed keypoint-level alignment (KP-Align) cannot be simply replaced by the naive classification-based solution (KP-CLS).
-1) Score Regression: Compared with Direct Regression-based (i.e., InternVideo2-MLP and USDL) and Multi-Task Learning-based (i.e., SwinBERT, PGMI) AQA methods, TechCoach achieves the best performance on score regression.
-Note that CoRe and TPT adopt a Contrastive Regression framework that utilizes 10 extra exemplars during inference, which brings stronger score regression ability.
-2) Coaching Commentary Generation: Although recently proposed Multimodal Large Language Models (MLLMs) show great visual context understanding and reasoning abilities, most of them [41, 14, 16] still struggle to handle realistic coaching scenarios (under-perform the task-specific models).
-Notably, the best among the evaluated MLLMs is the InternVL2 [6], which shows comparable performance with the task-specific models on LLM-based metrics.
-While InternVL2 is able to mention similar technical details as the ground truth (with a 47.76 GPT-M score), it falls short in providing precise guiding opinions (only 34.01 on the GPT-Q score).
-Among all the compared methods, our TechCoach achieves the best performance among all the metrics on coaching commentary generation.
-Variants Score Commentary
+- More Comparison Results.
+In Appendix, we also provide deeper explorations and analysis on various aspects, including (i) User study, (ii) Adding general TechPoints into the prompts, (iii) Abnormal performance of InternVL2 on traditional language metrics, (iv) Fine-tuned MLLM.
+{NiceTabular} l|cc|cccc Variants Score   Commentary
 […]
-70.83 4.49 36.25 13.34 15.88 66.36
-71.96 4.47 35.99 13.24 15.95 66.51
-72.00 4.39 36.43 12.57 16.07 66.72
-72.46 4.35 37.06 14.62 16.39 66.89
-Table 5: Ablation studies on the impact of Keypoint Features Augmentation. K.P.Text: Keypoint text features. K.P.Type: keypoint-type prompts. Quality: quality-related prompts.
-Variants Score Commentary
+w/o TP-Align 70.23 4.77 36.42 14.26 16.09 65.72
+TP-CLS 72.13 4.45 36.78 13.69 16.19 64.69
+TP-Align 72.46 4.35 37.06 14.62 16.39 66.89
+Table 4: Ablation studies on the alternative solutions of TechPoint-level alignment.
+The results show that the proposed TechPoint-level alignment (TP-Align) cannot be simply replaced by the naive classification-based solution (TP-CLS).
+{NiceTabular} ccc|cc|cccc Variants Score Commentary
 […]
-Ego-Only 69.31 4.75 36.87 13.79 16.16 66.13
-Exo-Only 68.55 4.84 37.31 12.71 16.20 66.23
-Ego+Exo 72.46 4.35 37.06 14.62 16.39 66.89
-Table 6: Comparison with Ego- and Exo-only variants.
-By collaborating on both Ego and Exo videos, TechCoach achieves stronger performance, showcasing the complementary role across different views for descriptive action coaching.
-Figure 5: Qualitative results on the generated coaching commentary.
+✓ 70.83 4.49 36.25 13.34 15.88 66.36
+✓ 71.96 4.47 35.99 13.24 15.95 66.51
+✓ ✓ 72.00 4.39 36.43 12.57 16.07 66.72
+✓ ✓ ✓ 72.46 4.35 37.06 14.62 16.39 66.89
+Table 5: Ablation studies on the impact of TechPoint Features Augmentation. T.P.Text: TechPoint text features. T.P.Type: TechPoint-type prompts. Quality: quality-related prompts.
+Refer to caption Figure 6: Qualitative results on the generated coaching commentary.
 Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos.
@@
 5.4 Ablation Studies
-- Is keypoint-level alignment a necessary?
+- Is TechPoint-level alignment necessary?
 We study this question by gradually ablating the training losses illustrated in Eq.
 3 from our full model.
-The results are shown in Tab.
-3.
+The results are shown in Sec.
+5.3.
 By comparing Row 3 with Row 1-2, it can be clearly observed that the model achieves unstable performance when simply combining the $L_{MSE}$ and $L_{MTM}$.
-This may caused by the gap between the score regression and commentary generation tasks.
-However, when further combining the keypoint-level alignment loss (Row 4), we can see stable performance improvement on all the compared metrics.
-Hence, we can conclude that keypoint-level alignment is necessary for the trade-off of score regression and coaching commentary generation.
-- Does keypoint-level alignment equal to binary classification?
-The main target of the keypoint-level alignment (KP-Align) is to ensure the Keypoint-related Quality Embeddings (see Fig.
+However, when further combining the TechPoint-level alignment loss ($L_{Align}$), we observe stable performance improvement on all the compared metrics (Row 4).
+Results in Fig.
+5 further illustrate how $L_{Align}$ impacts the performance.
+Starting from SwinBERT, after adding all our designs except $L_{Align}$, model performs much better on mentioning the technical details (3.04 improvement on GPT-M Score) due to the TechPoints inputs, but does not obtain similar improvement on perceiving action quality (only 1.45 improvement on GPT-Q Score).
+After further adding $L_{Align}$, much stronger improvement is observed on GPT-Q Score.
+All these results indicate the TechPoint-level alignment is necessary and will bring stronger action quality perception ability.
+- Does TechPoint-level alignment equal to binary classification?
+The main target of the TechPoint-level alignment (TP-Align) is to ensure the TechPoint-related Quality Embeddings (see Fig.
 3 and Sec.
 4.4 for more details) carry the quality-related information.
-To achieve this, a simple solution is to perform binary classifications over the Keypoint-related Quality Embeddings (i.e., classify whether the execution shows strengths or weaknesses on each keypoint), which we note as KP-CLS.
-To go deeper, we conduct a study on the influence of such an alternative solution.
-As shown in Tab.
-4, by comparing Row 2 with Row 1, we observe such a solution strengthens the quality perception of the model, resulting in performance improvement on score regression metrics.
-However, without explicit alignment with the keypoint-level coaching commentary, Bi-CLS falls short in understanding deeper relationships between the action execution and keypoints, resulting in a clear performance gap between our keypoint-level alignment, especially on commentary generation.
-- Impacts of the Keypoint Query Augmentation.
+To achieve this, a simple solution is to perform binary classifications over the TechPoint-related Quality Embeddings (i.e., classify whether the execution shows strengths or weaknesses on each TechPoint), which we note as TP-CLS.
+To dive deeper, we conduct a study on the influence of such an alternative solution.
+As shown in Sec.
+5.3, by comparing Row 2 with Row 1, we observe such a solution strengthens the quality perception of the model, resulting in performance improvement on score regression metrics.
+However, without explicit alignment with the TechPoint-level coaching commentary, Bi-CLS falls short in understanding deeper relationships between the action execution and TechPoints, resulting in a clear performance gap between our TechPoint-level alignment, especially on commentary generation.
+- Impacts of the TechPoint Query Augmentation.
 As discussed in Sec.
-4.4, we construct keypoint queries by augmenting the keypoint text features (K.P.Text) with keypoint-type prompts (K.P.Type) and quality-related prompts (Quality).
-As shown in Tab.
-5, we gradually ablate the components from the augmentation process, and the performance deterioration shows the contribution of each component.
-- Impacts of videos from various views.
-As discussed in Sec.
-5.2, we train TechCoach with both Egocentric and Exocentric videos.
-Here we conduct an experiment on the influence of the training videos from different views.
-As shown in Tab.
-6, Ego-Only and Exo-Only variants showcase comparable performance.
-By combining videos from different views, TechCoach achieves stronger performance, which shows that egocentric and exocentric viewpoints are complementary for action coaching.
+4.4, we construct TechPoint queries by augmenting the TechPoint text features (T.P.Text) with TechPoint-type prompts (T.P.Type) and quality-related prompts (Quality).
+In Sec.
+5.3, we gradually ablate the components from the augmentation, and the performance deterioration shows the contribution of each component.
+- More ablation studies.
+We also provide analysis on other aspects like (i) Impacts of various types of TechPoint-level alignment loss; (ii) Influence of the training videos from different views.
+Please refer to the Appendix for more details.
 5.5 Qualitative Results
@@
 As shown in Fig.
-5, TechCoach is able to understand various technical keypoints of different actions and provide detailed commentary on what is done well (e.g., “the extension of the shooting arm and wrist action were also commendable”) and what can be improved (e.g., “improvements are needed in core engagement to prevent being pulled from the wall”) from the given action videos.
+6, TechCoach is able to understand various technical TechPoints of different actions and provide detailed commentary on what is done well (e.g., “the extension of the shooting arm and wrist action were also commendable”) and what can be improved (e.g., “improvements are needed in core engagement to prevent being pulled from the wall”) from the given action videos.
 More qualitative results are shown in the Appendix.
 6 Conclusion
-In this paper, we investigate Descriptive Action Coaching, a novel task that focuses on providing coaching feedback on what is done well and what can be improved from an action execution.
-To achieve this, we propose TechCoach, a new framework that incorporates keypoint-level reasoning into the action coaching process, just as human coaches do.
-To support this task, we devise an automated pipeline for constructing the EE4D-DAC dataset which features clean and detailed coaching commentary on both keypoint and instance levels.
-Extensive experiments highlight the strong performance of the proposed TechCoach and demonstrate the effectiveness and the necessity of the proposed keypoint-level reasoning.
+We investigate Descriptive Action Coaching, a novel task that aims to provide coaching feedback on what is done well and what can be improved from an action execution.
+To support this task, we develop an automated pipeline for constructing the EE4D-DescCoach dataset which features clean and detailed coaching commentary on both TechPoint and instance levels.
+The TechPoint-level commentary provides new supervision for incorporating TechPoint-level reasoning into the action coaching process, encouraging us to build TechCoach, a new framework empowered by a Context-aware TechPoint Reasoner.
+Besides demonstrating strong performance of the proposed TechCoach, extensive experiments also highlights the effectiveness and the necessity of the proposed Context-aware TechPoint Reasoner.
+In Appendix, we further discuss the limitations of our work and outline future research directions, including (i) Scaling up TechCoach, (ii) Hierarchical TechPoints, and (iii) Pose-Assisted DescCoach.
 We expect the proposed new task, dataset, and method provide a new perspective for advancing current AQA to a more explainable and practical scenario.
 Appendix A Appendix
-In this appendix, we provide more details about the evaluation metrics, implementations, and qualitative results in Appendix B.
-Subsequently, more details about the EE4D-DAC dataset are introduced in Appendix C, including examples of general technical keypoints, hierarchical coaching commentary, prompts used for LLM-based annotation collection, and other details.
+In this appendix, we provide more details about the evaluation metrics, implementations, experiment results and analysis in Appendix B.
+Subsequently, more details about the EE4D-DescCoach dataset are introduced in Appendix C, including examples of general technical points (TechPoints), hierarchical coaching commentary, prompts used for LLM-based annotation collection, and quality ensurance.
 Finally, we discuss the limitations and potential directions for future work in Appendix D.
@@
 Spearman’s rank correlation coefficient ($\rho$) [41] is the most widely used metric to evaluate models in the field of AQA.
-It measures the correlation between the predicted score series and the ground-truth score series, which is defined as follows:
-$\rho=\frac{\sum^{N}_{i=1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum^{N}_{i=1}(x_{i}-\bar{x})^{2}\sum^{N}_{i=1}(y_{i}-\bar{y})^{2}}},$ (4)
-where $x$ and $y$ are the rankings of the two series.
-Relative-L2 Distance (RL2) [50] evaluates the numerical differences between the predicted and ground-truth scores, which can be written as:
-$RL2=\frac{1}{N}\sum^{N}_{n=1}(\frac{|s_{n}-\hat{s}_{n}|}{s_{max}-s_{min}})^{2},$ (5)
-where $s_{n}$ and $\hat{s}_{n}$ indicate the ground-truth and predicted scores, respectively. $s_{max}$ and $s_{min}$ indicate the highest and lowest scores for an action instance.
+It measures the correlation between the predicted score series and the ground-truth score series, which is defined as follows: $\rho=\frac{\sum^{N}_{i=1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum^{N}_{i=1}(x_{i}-\bar{x})^{2}\sum^{N}_{i=1}(y_{i}-\bar{y})^{2}}},$ (4) where $x$ and $y$ are the rankings of the two series.
+Relative-L2 Distance (RL2) [54] evaluates the numerical differences between the predicted and ground-truth scores, which can be written as: $RL2=\frac{1}{N}\sum^{N}_{n=1}(\frac{|s_{n}-\hat{s}_{n}|}{s_{max}-s_{min}})^{2},$ (5) where $s_{n}$ and $\hat{s}_{n}$ indicate the ground-truth and predicted scores, respectively.
+$s_{max}$ and $s_{min}$ indicate the highest and lowest scores for an action instance.
 BLEU [28] quantifies the overlap between the generated sentences and the reference sentences by comparing n-grams (contiguous word sequences) shared between them.
@@
 In our work, we use the DeBERTa-xlarge [11] to extract token-level embeddings as it shows great correlation with human evaluation.
-Figure 6: The detailed prompts for LLM-based evaluation. Zoom in for best view.
-Figure 7: An overview of the Spatial-Temporal Context Enhancer. Here the $\oplus$ indicates the element-level adding operation.
+Refer to caption Figure 7: The detailed prompts for LLM-based evaluation.
+Zoom in for best view.
+Refer to caption Figure 8: An overview of the Spatial-Temporal Context Enhancer.
+Here the $\oplus$ indicates the element-level adding operation.
 GPT-M and GPT-Q are the LLM-based metrics proposed to evaluate whether the generated commentary shares the same technical details and the same coaching opinions as the ground-truth commentary.
-We employ the GPT-4o-mini to determine the number of both-mentioned technical details (i.e., keypoints) and the shared coaching opinions.
+We employ the GPT-4o-mini to determine the number of both-mentioned technical details and the shared coaching opinions.
 Detailed prompts are provided in Fig.
-6.
+7.
 After that, the GPT-M score is calculated as the averaged proportion of the both-mentioned technical details to those mentioned by ground truth across all testing instances.
 The GPT-Q score is calculated as the averaged proportion of the technical details sharing the same coaching opinions to the both-mentioned technical details.
+Refer to caption Figure 9: Qualitative results on the LLM-based evaluation.
+We take genertaedt and ground-truth commentary from Fig.
+5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis.
+TD: recognized technical details.
 B.2 More Implementation Details
@@
 An overview of this module is shown in Fig.
-7.
+8.
 Specifically, we use a linear projector to map the clip-level features $F_{clip}$ to the hidden dimension $D$, and feed them into $N_{st}$ Transformer blocks that include spatial- and temporal-attention layers along with the FFN layers.
@@
 For the visual backbone, we use the pretrained weights of InternVideo2-stage3 [44].
+We use the pretrained DeBERTa-xlarge [11] as the text encoder.
 Following [19, 59], we employ AdamW [21] as the optimizer, and utilize a learning rate warm-up during the early 10% training steps followed by linear decay.
 $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 and 0.3, respectively.
-Due to the characteristics of EgoExo4D [10] and the annotation pipeline of EE4D-DAC, it is not possible to ensure each instance includes rounded commentary on every possible technical keypoints (e.g., in some cases the coaching commentary would not mention the execution about the head & eyes).
-Therefore, we make some adjustments on the prompts or instructions for the compared models to guide the models to focus on those technical keypoints mentioned in the ground-truth Commentary.
-For TechCoach, we employ zero-masking on the keypoint-related quality embeddings corresponded to a specific keypoint only if both the strength and weakness on this keypoint are unmentioned in the ground truth.
+Due to the characteristics of EgoExo4D [10] and the annotation pipeline of EE4D-DescCoach, it is not possible to ensure each instance includes rounded commentary on every possible TechPoints (e.g., in some cases the coaching commentary would not mention the execution about the head & eyes).
+Therefore, we make some adjustments on the prompts or instructions for the compared models to guide the models to focus on those TechPoints mentioned in the ground-truth Commentary.
+For TechCoach, we employ zero-masking on the TechPoint-related quality embeddings corresponded to a specific TechPoint only if both the strength and weakness on this TechPoint are unmentioned in the ground truth.
 Note that such a design does NOT leak the action quality information, and similar adjustments are applied on the compared models [59, 18, 15, 44, 6] for fair comparisons.
@@
 Besides, we modify the input templates for the text generator to include the score information and the technical dimensions that should be focused on.
-Specifically, the template is adjusted as follows:
-The player is practicing <Action_Task>, and gets a score of <Predicted_Score>. Pay attention to the execution of <Technical_Dimensions>”,
-where <Action_Task> is the action task, <Predicted_Score> indicates the predicted score by the model, and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground truth.
+Specifically, the template is adjusted as follows: The player is practicing <Action_Task>, and gets a score of <Predicted_Score>.
+Pay attention to the execution of <Technical_Dimensions>”, where <Action_Task> is the action task, <Predicted_Score> indicates the predicted score by the model, and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded TechPoints are mentioned in the ground truth.
 - Details about the Compared MLLMs [18, 15, 44, 6].
 We follow the official guidelines to perform zero-shot evaluation on the compared MLLMs.
-A general language instruction for these methods is as follows:
-Suppose you are an <Action_Task>coach.
+A general language instruction for these methods is as follows: Suppose you are an <Action_Task>coach.
 Watch the given video, and answer a paragraph of commentary about what is done well in the execution and what can be improved.
 Your answer should be like: <Answer_Example>.
-Pay attention to the execution of <Technical_Dimensions>”,
-where <Action_Task> is the action task (e.g., Reverse Layup), <Answer_Example> is an example of the answer (selected from training set), and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground-truth commentary.
-B.3 More Qualitative Results
+Pay attention to the execution of <Technical_Dimensions>”, where <Action_Task> is the action task (e.g., Reverse Layup), <Answer_Example> is an example of the answer (selected from training set), and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded TechPoints are mentioned in the ground-truth commentary.
+In Sec.
+B.3, we also explore building a DescCoach-specific MLLM.
+Specifically, we take VideoChat2 [15] as an example, and fine-tune it with LoRA [12] while keep the other parameters frozen on our EE4D-DescCoach dataset.
+Refer to caption Figure 10: User study.
+Our TechCoach wins the compared baselines frequently in human evaluation.
+Refer to caption Figure 11: Qualitative results on the LLM-based evaluation.
+We take genertaedt and ground-truth commentary from Fig.
+5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis
+{NiceTabular} l|cccccc Methods Commentary
+[…]
+VideoChat2-7B [15] 23.18 2.98 11.59 61.37 14.09 22.28
+VideoChat2-7B-FT [15] 28.74 4.88 13.27 62.98 36.32 23.62
+TechCoach(Ours) 37.06 14.62 16.39 66.89 50.44 38.15
+Table 6: Comparison with fine-tuned MLLM.
+While finetuning VideoChat2 on our EE4D-DescCoach dataset bring stable improvement on all metrics, the modest gain on GPT-Q score indicates that the model still struggle to obtain strong action quality perceiving ability.
+FT: finetuned on our EE4D-DescCoach dataset.
+{NiceTabular} l|cc|cccc Variants Score   Commentary
+[…]
+InfoNCE-Align 71.38 4.60 36.76 13.72 16.26 65.78
+MarginRanking-Align 71.67 4.36 36.96 14.29 16.35 66.45
+L2-Align 72.46 4.35 37.06 14.62 16.39 66.89
+Table 7: Comparison with variants with other types of feature alignment losses. TechCoach achieves stronger performance with simple yet effective L2-distance-based alignment loss.
+{NiceTabular} l|cc|cccc Variants Score   Commentary
+[…]
+Ego-Only 69.31 4.75 36.87 13.79 16.16 66.13
+Exo-Only 68.55 4.84 37.31 12.71 16.20 66.23
+Ego+Exo 72.46 4.35 37.06 14.62 16.39 66.89
+Table 8: Comparison with Ego- and Exo-only variants.
+By collaborating on both Ego and Exo videos, TechCoach achieves stronger performance, showcasing the complementary role across different views for descriptive action coaching.
+B.3 More Comparison Results and Analysis
+- User study.
+To further compare the performance across different models, we conducted a comprehensive user study comparing our TechCoach against three baselines (i.e., PGMI[59], SwinBERT[19], and InternVL2-8B[6]).
+We randomly select 50 instances from evaluation set and ask all four models to generate parallel outputs.
+We recruit five labelers (not involved in our project) and each of them is assigned instances through random sampling to eliminate evaluation bias.
+We provide similar assessment instructions as those for LLM evaluation (see Fig.
+7), and ask the labelers performed blind assessments by ranking the four outputs per instance based on semantic alignment with ground-truth (GT) text, allowing tied rankings (e.g., TechCoach=PGMI>InternVL2-8B>SwinBERT).
+After that, we computed pairwise comparison metrics: (1) Win Rate: Percentage of instances where TechCoach outperforms the baseline; (2) Tie Rate: Percentage of instances where there is a tie between TechCoach and the baseline; (3) Loss Rate: Percentage of instances where the baseline model prevail.
+The user study results are shown in Fig.
+10.
+Compared with other baselines, our TechCoach achieves clearly higher Win Rate, further illustrating the generated text of TechCoach can effectively catch the semantic meaning mentioned in Groun-Truth and align with human preference.
+- Adding general TechPoints into the prompts for MLLMs.
+As discussed in Sec.
+4.2, our TechCoach inferences with assumption that general TechPoints (GTP) are available.
+To evaluate the influence of GTP for general MLLMs, we try to add general TechPoints into the prompts for MLLMs.
+As shown in Fig.
+11, adding the (GTP) into the prompts brings unconsistent performance effects across different MLLMs (i.e., InternVL2-8B gets improvement but VideoChat2-7B and InternVideo2 gets opposite effects).
+These results suggest that a more effective approach to integrating TechPoints into the action assessment process is necessary.
+- Fine-tuning MLLM on EE4D-DescCoach dataset.
+As another branch of exploration, we try to fine-tune MLLM on our EE4D-DescCoach dataset.
+We take VideoChat2 [15] as an example, and fine-tune it with LoRA [12].
+The results are shown in Sec.
+B.2.
+While finetuning VideoChat2 on our EE4D-DescCoach dataset brings stable improvement on all metrics, the modest gain on GPT-Q score indicates that the model still struggle to obtain strong action quality perceiving ability.
+- Why does InternVL2 achieve abnormal performance on traditional metrics?
+In Tab.
+2 of main paper, InternVL2 [6] family models obtain abnormal performance on traditional n-gram based metrics (i.e., BLEU, METEOR, CIDEr) and feature similarity-based metric (BERTScore).
+By going deeper to the outputs of InternVL2, we find the reasons behind it could be: even if an answer sample is provided in the prompt, InternVL2 tends to provide answers in a fixed, list-like format.
+For example: ## Execution Done Well: 1.
+**Starting Position**: The player demonstrates a strong starting position with … 2.
+**Ball Handling**: The dribbling technique appears consistent … The formats of answers and ground-truth commenatry are highly inconsistent, which makes the traditional metrics fail to effectively evaluate the model performance.
+B.4 More Ablation Studies
+- Impacts of various types of TechPoint-level alignment loss.
+As discussed in Sec.
+4.4, L2 distance is used to measure the distance between TechPoint-related quality embeddings and TechPoint-level commentary features.
+We further investigate the imapcts of alignment loss by introducing two other kinds of alignment losses, i.e., InfoNCE and MarginRanking.
+As shown in LABEL:{tab:ablation_alignment_loss}, TechCoach achieves better performance with the L2 distance-based alignment loss.
+- Impacts of videos from various views.
+As discussed in Sec.
+5.2, we train TechCoach with both Egocentric and Exocentric videos.
+Here we conduct an experiment on the influence of the training videos from different views.
+As shown in Sec.
+B.2, Ego-Only and Exo-Only variants showcase comparable performance.
+By combining videos from different views, TechCoach achieves stronger performance, which shows that egocentric and exocentric viewpoints are complementary for action coaching.
+B.5 More Qualitative Results
+- LLM can provide a reliable evaluation of generated text.
+To highlight this, we take an example from Fig.
+5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis, which are shown in Fig.
+9.
+- TechCoach is able to perceive the action quality.
 As a supplement to Fig.
 5 in the main paper, we present additional qualitative results on the generated coaching commentary in Fig.
-8, further illustrating the effectiveness of our TechCoach.
+12, further illustrating the effectiveness of our TechCoach.
 Moreover, Fig.
 13 demonstrates that our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details on various action executions.
-Appendix C More Details about EE4D-DAC Dataset
-In this section, we show the examples of the obtained general technical keypoints, hierarchical coaching commentary, prompts for annotation collection, and other details.
-General Technical Keypoints.
+Appendix C More Details about EE4D-DescCoach
+In this section, we show the examples of the obtained general TechPoints, hierarchical coaching commentary, prompts for annotation collection, and other details.
+- General TechPoints.
 Completing a skilled physical action typically requires coordination across multiple body parts and objects.
-Thus, we define technical keypoints as language instructions spanning seven technical dimensions: 1) Head & Eyes; 2) Torso & Core; 3) Arms & Elbows; 4) Wrists & Hands; 5) Legs & Knees; 6) Feet & Heels; and 7) Human-Object Interaction.
-Tab.
-7, Tab.
-8 and Tab.
-9 provide examples of the general technical keypoints for specific action tasks.
+Thus, we define TechPoints as language instructions spanning seven technical dimensions: (1) Head & Eyes; (2) Torso & Core; (3) Arms & Elbows; (4) Wrists & Hands; (5) Legs & Knees; (6) Feet & Heels; and (7) Human-Object Interaction.
+Appendix D, Appendix D and Appendix D provide examples of the general TechPoints for specific action tasks.
 - Hierarchical Coaching Commentary.
-Our EE4D-DAC dataset features hierarchical coaching commentary on both keypoint level and instance level.
+Our EE4D-DescCoach dataset features hierarchical coaching commentary on both TechPoint level and instance level.
 Fig.
@@
 As described in Sec.
-3.2 of the main paper, we use LLM (GPT-4o) to generate general technical keypoints and hierarchical coaching commentary.
+3.2 of the main paper, we use LLM (GPT-4o) to generate general TechPoints and hierarchical coaching commentary.
 Additionally, as mentioned in Sec.
@@
 The detailed prompts are illustrated in Fig.
-12.
-Manual Check on Annotations. After obtaining the annotations with LLM, we manually check the evaluation set to ensure the annotation quality.
+16.
+- Commentary length and vocabulary size.
+As shown in Appendix C, the vocabulary size of EE4D-DescCoach dataset is 7297.
+The TechPoint-level and Instance-level commentary spans average lengths of 15.8 and 68.9, respectively.
+[…]
+TechPoint-level 15.8$\pm$4.33 7297
+Instance-level 68.9$\pm$14.1 7297
+Table 9: Lengths and vocabulary size of EE4D-DescCoach dataset.
+Manual Check on Annotations.
+After obtaining the annotations with LLM, we manually check the evaluation set to ensure the annotation quality with the help of self-developed visualization tool.
+We filter the misaligned instances or prompt the LLM to re-generate the annotation until it meets our requirements.
 Appendix D Limitations and Future Work
-Although experiments showcase the effectiveness of the proposed TechCoach, some limitations remain and warrant further exploration.
-Hierarchical Technical Keypoints.
-In this work, for a given action task, we assume that each body part (or the dimension of human-object interaction) is associated with one paragraph of technical keypoints.
-While this assumption provides a general framework for the initial exploration of technical keypoint-aware DAC, technical keypoints could be further constructed hierarchically.
+- Scaling up TechCoach.
+Our TechCoach is built with a relatively small text generator (136M).
+We conduct fair comparisons with task-specific models training on our dataset with the same text generator and visual backbone (InternVideo2), and provide details analysis to show the contribution of the proposed Context-aware TechPoint Reasoner.
+However, the small text generator with well large-scale pretraining may bring shortcomings on more difficult scenarios such as out-of-domain instances.
+Note that action coaching in OOD scenario is still an open problem in the related field.
+Due to the resource and computation constraint, we currently not able to integrate an LLM (7B or larger) with our TechCoach and train the Context-aware TechPoint Reasoner from scratch.
+We believe scaling up TechCoach, and conducting large-scale pre- and post-training will further enhance the DescCoach performance in various scenarios.
+We leave them for future work.
+- Hierarchical TechPoints.
+In this work, for a given action task, we assume that each body part (or the dimension of human-object interaction) is associated with one paragraph of TechPoints.
+While this assumption provides a general framework for the initial exploration of TechPoint-aware DescCoach, TechPoints could be further constructed hierarchically.
 For instance, when performing Mid-range Shooting, the shooting hand and guide hand play distinct roles in determining shooting accuracy and stability.
-Separating the technical keypoint for high-level body parts (e.g., wrists & hands) into finer sub-parts (e.g., shooting hand and guide hand) could further guide the model to focus on fine-grained and specific aspects in an action execution.
-Pose-Assisted DAC.
-In this work, we focus on video-based DAC.
+Separating the TechPoint for high-level body parts (e.g., wrists & hands) into finer sub-parts (e.g., shooting hand and guide hand) could further guide the model to focus on fine-grained and specific aspects in an action execution.
+- Pose-Assisted DescCoach.
+In this work, we focus on video-based DescCoach.
 While the Ego-Exo input enriches the visual context, there are cases where the egocentric video lacks sufficient visual information, and the human subject in the exocentric video appears relatively small.
 A valuable auxiliary modality to address this limitation is body pose.
-Since the original EgoExo4D [10] dataset provides frame-aligned 3D body pose data, incorporating pose sequences into the DAC process represents a promising direction for future research.
-Figure 8: More qualitative results on the generated coaching commentary.
+Since the original EgoExo4D [10] dataset provides frame-aligned 3D body pose data, incorporating pose sequences into the DescCoach process represents a promising direction for future research.
+Refer to caption Figure 12: More qualitative results on the generated coaching commentary.
 Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos.
 Correctly matched parts are highlighted in colors.
-Figure 9: More qualitative results on the generated coaching commentary.
+Refer to caption Figure 13: More qualitative results on the generated coaching commentary.
 Our TechCoach is capable of perceiving the action quality (i.e., Strengths and Weaknesses) of the same technical details (e.g., bending the knee) across various action executions.
@@
 Some frames are zoomed in for better visibility.
-Technical Dimensions General Technical Keypoints
+{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
 Head & Eyes The player should keep the head up and eyes focused on the basket to maintain good court awareness and accuracy.
@@
 Human-Object Interaction The player should practice alternating layups from one side of the basket to the other, using the backboard as a guide to ensure the ball goes into the basket consistently.
-Table 7: General Technical Keypoints for Basketball Drills - Mikan Layup.
-Technical Dimensions General Technical Keypoints
+Table 10: General Technical Points for Basketball Drills - Mikan Layup.
+{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
 Head & Eyes The climber should keep the head up and eyes focused on the next hold. Always plan your next few moves in advance and be aware of your surroundings to maintain good spatial awareness.
@@
 Use both hands and feet to distribute weight evenly and move smoothly from one hold to the next.
-Table 8: General Technical Keypoints for Rock Climbing.
-Technical Dimensions General Technical Keypoints
+Table 11: General Technical Points for Rock Climbing.
+{NiceTabular} p2.8cm|p11cm Technical Dimensions General Technical Points
 Head & Eyes The player should keep the head up and eyes focused on the ball at all times. This helps in tracking the ball’s movement and anticipating its next position.
@@
 Each touch should be calculated to keep the ball at a manageable height and within close proximity to the body.
-Table 9: General Technical Keypoints for Soccer Drills - Juggling.
-Figure 10: An example for hierarchical coaching commentary. Zoom in for best view.
-Figure 11: An example for hierarchical coaching commentary. Zoom in for best view.
-Figure 12: The detailed prompts for generating General Technical Keypoints (Upper), Hierarchical Coaching Commentary (Middle), and Coach-Scores (Lower).
+Table 12: General Technical Points for Soccer Drills - Juggling.
+Refer to caption Figure 14: An example for hierarchical coaching commentary. Zoom in for best view.
+Refer to caption Figure 15: An example for hierarchical coaching commentary. Zoom in for best view.
+Refer to caption Figure 16: The detailed prompts for generating General TechPoints (Upper), Hierarchical Coaching Commentary (Middle), and Coach-Scores (Lower).
 “<Action_Domain>”: The action domain of the instance (e.g., Basketball, Soccer, Roch Climbing).
 “<Action_Domain>”: The action task of the instance (e.g., Reverse Layup).
-“<General_Technical_Keypoints>”: The generated general technical keypoints for a specific action task.
+“<General_TechPoints>”: The generated general TechPoints for a specific action task.
 “<Original_Expert_Commentary>”: The original expert commentary on the instance given by EgoExo4D [10].
 Zoom in for best view.
+(a) Comparison with existing methods on score regression.